A Bibliography of Publications in IEEE Computer: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

13 November 2025 Version 2.20

Title word cross-reference

3 [321].

-D [321].

1 [392, 542]. 11 [45]. 1980s [414]. 1987 [411]. 1990s [83].

2 [249]. 2D [548].

3 [322, 549]. 3-D [322, 549]. 38 [360]. 3D [401, 548].

432 [277]. 4th [182].

5 [495].

754 [26]. 79 [32].

’84 [90, 117, 185, 289, 313]. 85 [149].

Abingdon [541]. Abstractions [343]. Access [449, 465]. accidents [383]. According [458]. Accreditation [171, 172]. Accuracy [511]. ACM [185, 439]. Acquisition [548]. across [504]. Ada [31, 123, 357, 365, 505]. Adaptable [17, 202]. Adaptive [397, 450, 454, 536]. Added [331]. Addressable [544]. ADMS [371]. ads [108, 131, 159, 190, 214, 233, 267, 293]. Advanced [119, 317, 498, 528]. aerodynamics [555]. Against [177]. Aided [424]. Alamos [552]. Algebra [393]. Algorithm [124, 246, 579]. Algorithmic [351, 509, 534]. Algorithms [8, 49, 323, 327, 375, 396, 483]. Allocation [19]. Alternative [175]. Alternatives [544]. Ambler [532]. Amdahl [482]. Ames [555]. Analog [452]. Analysis [25, 41, 77, 341].

Animating [340]. Announcements [101, 102, 127, 128, 156, 157, 180, 181, 207, 208, 229, 230, 259, 260, 284, 285, 422, 423]. Annual [184, 210]. Appendum [485]. Application [94, 467]. Applications [26, 67, 73, 103, 121, 123, 129, 155, 189, 206, 231, 261, 275, 286, 315, 328, 474, 554]. Applied [124]. Applying [64]. Appreciation [457, 485]. Apprentice [486]. Approach [66, 72, 79, 200, 312, 358, 564, 567]. Approaches [3, 142, 468]. April [485]. Arbitrary [483]. Architectural [148]. Architecture [14, 34, 73, 95, 274, 277, 334, 347, 380, 381, 386, 448, 474, 477, 482, 500]. Architecture/Compiler [386]. Architectures [17, 50, 71, 352, 376, 378, 489]. area [12]. Arithmetic [2, 23, 24, 26, 51, 201]. Armstrong [467]. Array [34–36, 391, 394, 396, 397, 483]. Arrays [246, 353, 375, 390, 395, 399, 400, 461]. ART [454]. Artificial [197, 302, 312, 376, 378]. assembly [514]. Assistant [361, 522]. Associate [170]. Associative [544]. Assurance [177]. Assuring [560]. Astronautics [542]. Augmented [388]. Authority [449]. automated [514]. Automatic [62, 475, 520]. Automating [310]. Automation [30, 508, 513–516]. automation-robotics [516]. Autonomous [280, 310, 531–533]. Availability [276]. Avalon [493]. Avalon/C [493].

Back [3, 4]. Back-End [3, 4]. Balancing [58]. band [12]. Base [39, 40, 64, 380, 433]. Based [6, 58, 81, 250, 303, 346, 359, 374, 401– 406, 519, 577]. batch [514]. Beamforming [397]. Behavior [80, 251]. benchmark [541]. Beyond [342]. Bibliography [456]. Bidirectional [459]. Binary [23, 452, 458]. Birth [497]. Book [109, 135, 163, 192, 215, 236, 268, 296, 345, 429]. Bookshelf [110, 136, 164, 193, 216, 237, 269, 297]. Box [466]. Box-Structured [466]. Branch [88].

Broadband [568]. Broadcast [8]. Browsing [358, 443]. Buffer [45, 88]. Building [81, 308, 384, 539]. bullet [383]. Bus [16]. Business [30].

C [493]. C2 [496]. Cabinet [465]. caches [491]. CAD [323, 379, 401, 402, 405, 525]. CAD-based [401]. CAE [184, 210]. calculating [9]. Calendar [106, 132, 160, 187, 211, 234, 265, 294, 426]. Call [107, 133, 161, 188, 212, 235, 253, 266, 295, 311, 427]. CAM [323]. Camera [330]. Can [225]. Capacity [337, 360]. Career [428]. Case [124, 415, 432, 491]. CCD [460]. Cellular [95, 398]. Center [554, 555]. Chaining [559]. Channel [97, 151, 203, 226, 252]. Characterization [472]. Charlotte [561]. Checking [7]. Chief [140, 410]. Chinese [324]. Chinese/Kanji [324]. Chip [32, 459]. Chips [1]. Circuit [567]. Circuits [7, 459, 460, 520]. Classified [108, 131, 159, 190, 214, 233, 267, 293]. Cm* [68]. Code [247, 504]. Codes [7]. coding [580]. Coherence [445]. Collaboration [254]. Collapsed [521]. Collapsed-Fault [521]. College [173]. Color [78]. Combatting [176]. Combinatorial [335]. Combining [526]. Commercial [275, 384, 449]. Committee [96]. Communication [12, 66, 221]. Communications [149]. Comparison [458]. Compcon [90, 117, 288]. Compiler [11, 386, 505]. Compiler-Compiler [11]. Compiling [247]. Complete [370, 583]. Complex [415]. Complexity [342]. Components [355]. Composing [502]. Compression [220, 579, 580, 583]. Compsac [313]. Computation [15, 49, 463, 511]. computational [555]. Computations [564]. Computer [22, 38, 55, 58, 63, 66, 71, 74, 86, 89, 94, 96, 98, 105, 113, 115, 120, 123, 130, 139– 141, 144, 147, 150, 158, 167, 169– 172, 174, 183, 209, 219, 223, 232, 240, 255, 263, 272, 277, 280, 290, 291, 300, 308, 316, 318, 320, 330, 334, 346, 376, 380, 386, 392, 409, 411, 419, 424, 431, 439, 489, 498, 507, 571, 575, 577]. Computer-Aided [424]. Computer-Based [346]. Computer-Controlled [120]. Computers [121, 146, 147, 149, 168, 173, 255, 274, 290, 315, 342, 414, 473, 478, 545]. Computing [12, 35, 48, 51, 182, 184, 210, 254, 273, 367, 439, 452, 455, 481, 506, 571]. Concept [390, 391, 400, 444]. Concepts [281, 328, 479, 528]. Concurrent [51, 326, 473, 564]. Conference [121, 123, 134, 147, 182, 255, 257, 290, 309, 315, 411, 424, 480]. Conferences [425]. Conferencing [346]. Configuration [331, 566]. Configurations [356]. Confluence [149]. CONLAN [328]. Connection [474]. Connectionist [378]. Considerations [6, 61, 374, 499]. Constrained [537]. Constraints [578]. Construction [322, 444]. Content [544]. Content-Addressable [544]. Contention [487]. contents [85, 112, 138, 162, 166, 186, 195, 213, 218, 239, 262, 271, 287, 299, 408]. Control [114–116, 118, 119, 146, 280, 362, 403, 406, 449]. Controlled [120]. Controller [397]. Controlling [536]. Controls [314]. Controversy [342]. Convex [496]. Cooperation [152]. Core [439]. Correcting [7]. Correctness [470, 488]. Coupled [447]. cover [111, 137, 165, 194, 217, 238, 270, 298, 407]. Coverage [521]. Covered [370]. Cray [249, 497, 559]. CRIB [150]. Criteria [458]. Critical [94]. Cross [541]. Crossbar [389]. Culture [140]. Current [338, 347, 459, 470, 513–516]. Current-Mode [459]. Curriculum [562]. Curved [69]. Cut [331]. Cydra [495].

D [321, 322, 549]. Dallas [411]. DARPA [74, 349]. Data [18, 19, 39–42, 53– 56, 59, 64, 134, 146, 220, 248, 323, 324, 380, 400, 433, 468, 483, 547–550, 556, 574, 577, 579, 583]. data-base [40]. Data-Driven [400, 483].

Data/Knowledge [380, 433]. Database [71, 91, 202, 306, 307, 320, 354, 358, 573, 575]. Databases [344, 369, 373, 379, 432, 578]. Dataflow [52, 351]. Deadlock [13, 569]. Debugging [565]. Decade [149]. Decentralized [280]. Decision [312]. decisions [495]. Declarative [304, 564]. Decomposing [323]. defense [501]. Definitive [582]. Degree [170]. Denormalized [27]. departmental [495]. Depth [539]. Description [304, 329]. Design [1, 6, 45, 66, 88, 105, 116, 118, 124, 130, 147, 158, 183, 198, 202, 209, 232, 255, 263, 281, 290, 291, 308, 326, 343, 354, 357, 362, 363, 371, 374, 397, 416, 424, 434, 471, 490, 495, 496, 499, 522, 543, 567, 577]. Designing [276, 561]. Designs [519]. detection [569]. Determining [78]. Development [10, 222, 339, 435, 440, 464, 519, 523]. Developments [394]. Diagnostic [306, 307]. Digital [80, 200, 448]. direct [491]. direct-mapped [491]. Directed [306, 307]. Directions [14]. Disaster [177]. Discipline [439, 506, 507]. Disk [59, 337, 374]. Disk-Based [374]. Disks [465]. Display [547]. Distributed [8, 12, 19, 67, 89, 94, 115, 119, 175, 182, 221, 350, 357, 387, 492, 538, 552, 569]. distributions [9]. Diversity [281]. do [382]. Document [347, 440, 441]. domain [482]. Domesticating [366]. Down [322]. drawing [5]. Drawings [322]. Driven [400, 471, 483]. Dual [402]. Dynamic [484, 542].

Earth [550]. easy [5]. easy-to-use [5].

Editing [441]. Editor

[87, 114, 140, 168, 301, 302, 387, 410, 456]. Editor-in-Chief [140, 410]. Editors [84, 242, 273, 436, 573]. Education [168, 174, 184, 210, 507]. Effect [242]. Efficient [326, 370]. Electrical [184, 210, 431]. Electronic [76, 345, 515, 551]. Electronics [431]. Encryption [176]. End [3, 4]. Engineer [522]. Engineering [134, 150, 169, 184, 210, 224, 322, 344, 383, 496, 526, 546, 562]. Engineers [431]. Enhancement [464]. Enhancing [344]. Environment [28, 29, 31, 168, 222, 245, 319, 358, 364, 438, 444, 463, 522, 539]. Environments [123, 434, 435, 437, 563]. Errata [24]. Erratum [104, 264, 292]. Error [7]. Error-Correcting [7]. essence [383]. Essential [332, 396]. Ethernet [63]. Evaluating [356]. Evaluation [384]. Event [445]. Events [191]. Evolution [63, 524, 563, 571]. Examples [201]. Exception [512]. Exchange [388]. Execution [46, 246, 251]. Exercise [471]. Expanding [546]. Expansion [197]. Experience [81, 561]. Experimental [349, 556]. Experiments [62, 281]. Expert [303, 306, 307, 360, 362–364, 384, 479]. Exploration [532]. Expressing [327].

Facilitate [36]. Facility [463, 528, 561]. Facts [443]. Fall [117, 288, 411]. Fault [150, 273–275, 277–279, 281, 395, 447, 521]. Fault-finding [150]. Fault-Tolerant [273–275, 277–279, 447]. Federal [333]. Field [547]. Fifth [141, 334]. Fifth-Generation [141, 334]. File [62, 89, 380, 465]. Filtering [450]. Finding [150, 443]. First [121, 309, 315]. Fixed [375]. Fixed-Size [375]. Flames [548]. Flexible [67]. Flight [528]. Floating [2, 23–26]. Floating-Point [2, 23, 25]. Flow [18, 53–56, 64, 146, 248, 547]. Flows [548]. Fluid [547]. FolioPub [442]. Force [439]. Formal [416, 470]. Formalization [437]. Format [581]. Formats [347]. Fortran [36, 469]. Framework [198, 525, 534]. Free [60]. Friends [366]. Front [111, 137, 165, 194, 217, 238, 270, 298, 407]. Functional [79, 368]. Functionally [34]. Future [146, 191]. Fuzzy [479].

GaAs [386, 397]. Gateway [47]. Gateways [73, 74]. General [392, 499]. General-Purpose [392, 499]. Generating [521]. Generation [20, 141, 143, 334, 481, 518, 520]. Genesis [320]. Gentest [520]. geographic [41, 574]. GEOMATIC [321]. Geometric [511]. Good [560]. Government [254, 333]. Gradual [197]. Graduation [530]. Grain [248]. graph [5]. Graphic [321, 581]. Graphical [319, 338, 339, 492, 566]. Graphics [22, 316, 348, 477]. graphs [54]. Grasping [513]. Gridnet [175]. Grids [535]. Guest [84, 114, 168, 242, 273, 302, 387, 456, 573]. Guide [2, 24, 582]. Guidelines [170].

Handling [363, 512]. Hardware [80, 242, 325, 327, 329, 470, 522, 572]. Hardware-Software [242]. Hashing [484]. Help [225]. Heterogeneous [74, 91, 463]. High [142, 220, 337, 448, 478, 518]. High-Level [518]. High-Performance [220]. High-Speed [337, 448, 478]. High-Technology [142]. Highly [48, 51]. Holy [37]. Hyper [468]. Hyper-Semantic [468]. Hyperchannel [75, 199]. Hypertext [412, 443].

IBM [360, 469, 557, 559]. IC

[101, 127, 156, 180, 207, 229, 259, 284, 422]. ICs [458]. IEEE [26, 86, 96, 98, 113, 123, 139, 147, 149, 167, 169, 219, 240, 255, 272, 290, 300, 424]. IEEE-754 [26]. If [60]. Image [71, 573, 576]. impact [12]. Implementation [2, 24, 370, 390, 391, 400, 451]. implementations [328]. Improving [222, 413]. Inc. [431]. individual [580]. Industrial [114, 115, 119, 122]. Industry [145, 254]. Influence [563]. Infocom [149]. Infomart [411]. Information [41, 42, 105, 130, 158, 183, 209, 232, 263, 291, 348, 444, 490, 539, 574]. Inheritance [493]. Inspection [540]. inspections [488].

Institute [431, 530, 562]. Instruction [243, 244, 342, 415, 542]. Instruction-Level [243, 244]. Insurance [177]. Integrated [71, 434, 577]. Integrating [35, 354, 399]. Integration [118, 149]. Integrity [578]. Intel [277]. Intelligence [197, 302, 312, 376, 378]. Intelligent [47, 513, 531]. Interaction [242]. Interactive [184, 200, 210, 316, 319, 549]. Interchange [347, 348]. Interconnecting [74, 76, 91, 355]. Interconnection [43, 44, 199, 387, 388, 503]. Interface [302, 304, 343, 527]. Interfaces [305, 492, 502]. Interfacing [323]. Interlisp [29]. Intermedia [444]. Internal [484]. International [121, 134, 147, 182, 189, 255, 257, 290, 309, 315, 347, 424]. Internet [74]. internetworking [73]. Interoffice [350]. Interpreter [70]. Interpreting [33]. Intersociety [152]. InterViews [502]. Introduction [84, 114, 148, 168, 242, 273, 302, 387, 412, 456, 470, 472, 573]. investigate [488]. IS-A [77]. ISDN [568]. Isn’t [77]. ISR [575]. Issues [4, 14, 308, 332, 357, 505].

Japan [143, 144, 224]. Japanese [140–142, 145]. Job [310]. Job-Shop [310]. Joint [411]. Jukebox [337].

Kanji [324]. Kernel [538]. Key [176]. Kingdom [394]. Knowledge [79, 81, 82, 150, 303, 306, 307, 344, 362, 380, 433, 443, 526, 577]. Knowledge-Based [81, 303, 577]. Knowledge-Directed [306, 307]. KRYPTON [79].

Laboratory [552, 553, 556]. LANDSAT [41]. Language [10, 14, 30, 305, 326, 327, 329, 436, 563, 577]. Languages [40, 53, 56, 119]. Large [15, 49, 175, 248, 337, 353, 380, 432, 550]. Large-Capacity [337]. Large-Grain [248]. Large-Scale [15]. Lawrence [553]. learned [534]. Learning [168, 533, 537]. length [9].

Letters [87, 301]. Level [173, 243, 244, 510, 518]. Library [519]. Library-Based [519]. Linda [366]. Linear [393]. Links [75, 77, 199]. LISP [32, 381, 472]. List [196]. Literature [100, 126, 154, 179, 205, 228, 258, 283, 430]. Livermore [553]. Load [58]. Local [12, 58, 63, 199]. Localization [405]. Logic [146, 456, 457, 459, 461, 485].

Machine [247, 302, 334, 474]. Machines [56, 64, 381, 433, 531]. Magazine [139, 167]. Mail [76, 349, 350]. Maintenance [543]. Man [302]. Man-Machine [302]. Management [57, 71, 91, 122, 176, 312, 372, 442, 480, 527, 573, 577]. Managing [379]. Manipulating [42]. Manipulation [540]. Manipulators [406]. Manufacture [200]. Manufacturing [310, 405]. Map [574]. mapped [491]. Mapping [375, 396, 490]. Mass [60, 61, 336]. Master [562]. Matrix [51, 392]. Matrix-1 [392]. matures [507]. Maximum [469]. Measuring [9, 69]. Mechanical [322, 362, 514]. Mechanism [304]. Mechanisms [221]. Medical [549]. Member [431]. Memories [278, 279]. Memory [38, 57, 487, 544]. Message [241, 409, 410, 473]. Message-Passing [473]. Metadata [372]. Methodology [118, 202]. Methods [416, 545]. Microcomputer [16]. Microcomputers [189]. Microprocessor [6, 92, 93, 499]. Microprocessor-Based [6]. Microsystem [128, 157, 181, 208, 230, 260, 285, 423]. Microsystems [102]. Mid [149, 414]. Mid-1980s [414]. Mid-Decade [149]. Migration [62, 561]. Mini [382]. Mini-Super [382]. Misconceptions [481]. MML [92, 93]. Mobile [533, 535]. Mode [459]. Model [94, 169, 312, 336, 404–406, 451, 464, 539, 576]. Model-Based [404–406]. Modeling [243, 244, 246, 323, 379, 462, 468]. Modular [43]. Monitoring [46]. Motion [33, 452, 509]. Motions [510]. MP [249, 497, 559]. Multicomputer [356, 462]. Multicomputers [473]. Multilevel [73]. Multimedia [349, 374]. Multiple [92, 93, 440, 456, 457, 459–461, 482, 485]. multiple-domain [482]. Multiple-Microprocessor [92, 93]. Multiple-Valued [456, 457, 459–461, 485]. Multiplier [459]. Multiprocessing [335]. Multiprocessor [247, 249, 250, 332, 447, 467, 503]. Multiprocessors [351, 445, 487]. Multisensor [538]. Multistage [388]. Multitasking [249]. Multivalued [458]. Myths [475].

NASA [555]. National [480, 552–554]. Natural [305]. Natural-Language [305]. Nature [384]. Naval [556]. Navigation [533–535]. Nets [450]. Network [3, 58, 63, 66, 75, 175, 199, 446, 451, 453, 454, 529, 568, 581]. Networks [4, 21, 43, 44, 49, 74, 77, 223, 387– 389, 452, 455, 503]. Neural [450, 451, 453–455, 537]. News [419]. Next [143, 481]. Next-Generation [481]. No [383]. Nonlinear [537]. Null [397]. Numbers [27].

OBE [30]. Object [322, 374, 401, 525, 560, 576]. Object-Oriented [525, 560]. Objects [33, 373]. Occluded [404]. Occupancy [535]. October [411]. Office [30, 347]. Offs [45, 495]. Once [465]. Open [97, 151, 203, 226, 252]. Operating [57, 276, 317, 357, 446]. Operator [361]. Opinion [56]. Opportunities [428, 546]. Optical [59, 337, 374, 389, 433, 465, 568]. Optimization [461, 472, 504]. Optimum [448]. Ordering [445]. Organization [45]. Organizing [454]. Oriented [525, 560]. Overlapping [330]. Overview [11, 13, 116, 416, 486]. package [5]. Packaging [96]. Packet [43]. Paper [363]. Papers [107, 133, 161, 188, 212, 235, 253, 266, 295, 311, 427]. Para [368]. Para-Functional [368]. Paradigm [83]. Parallel [34, 48, 49, 245, 248, 254, 257, 333, 334, 352, 365, 367, 387, 393, 432, 448, 545]. Parallelism [366, 385]. PARET [462]. Part [78]. Partially [404]. Partitioning [375, 396]. Parts [330, 404, 405]. Passing [473]. Patrec [306, 307]. Pattern [318, 450, 453, 454]. PC [557]. peace [37]. Perception [535]. Performance [4, 220, 242, 250, 341, 351, 356, 374, 469, 503, 565]. Persistent [72]. Personal [173]. Perspective [31, 571]. Perspectives [15]. philosophies [495]. Pick [510]. Pick-and-Place [510]. PICT [319]. Pictorial [39–42, 578]. Picture [40]. Pipelining [559]. Piracy [176]. pivotal [501]. Place [510]. Placement [89]. Planetary [532]. Planning [360, 404, 509, 510, 514]. Plausibility [471]. Plausibility-Driven [471]. plea [37]. PNG [581, 582]. Point [2, 23–26]. Poker [245]. Portable [581]. Power [361]. Practical [55, 438]. Practices [224]. Precision [500]. Precollege [174]. Precompilation [36]. Predictable [538]. Prediction [88, 403]. Prediction-Based [403]. Preliminary [439]. President [409]. Price [331, 377]. Pride [363]. Principles [371]. Problem [13, 481]. Problems [94, 335, 511]. procedures [504]. Process [114, 115, 118, 119, 314, 561]. Processing [19, 36, 41, 51, 71, 248, 257, 324, 365, 376, 387, 398, 436, 447, 448, 517, 574, 575]. Processor [35, 243, 244, 246, 331, 333, 353, 393, 400, 472]. Processors [34, 391, 396, 414]. Product [177, 420]. Production [11, 122, 359]. Productivity [222, 413]. products [99, 125, 153, 178, 204, 227, 256, 282, 421]. Profession [171]. professional [105, 130, 158, 183, 209, 232, 263, 291]. Program [46, 54, 169, 172, 243, 244, 339, 416, 436, 488]. Programmable [461, 529]. Programmer [486]. Programming [28, 29, 72, 92, 93, 221, 245, 319, 338, 368, 385, 406, 434, 437, 438, 469, 475, 514, 566, 577]. Programs [36, 170, 333, 340, 560]. Project [11, 141, 328, 554]. PROLOG [326]. Properties [493]. Proposals [25]. Proposed [2, 16, 23, 24, 26]. Prospects [475]. Protocols [568]. Prototypes [527]. Prototyping [515, 523, 524, 526, 528, 529]. PSN [81]. Publication [442]. Purpose [1, 392, 499].

QBE [30]. QBE/OBE [30]. Quality [11]. Quantum [551]. Query [40, 517]. Query-Processing [517]. Querying [432]. queue [9].

R&D [142, 480]. RAM [544]. Random [465]. Random-Access [465]. Rapid [523, 524, 526, 528]. rate [580]. Reaching [22]. Reading [345]. Real [67, 119, 346, 448, 481, 538]. Real-Time [67, 119, 346, 448, 481, 538]. Recognition [318, 330, 450, 453, 454, 576]. Recognizing [404]. Recommendations [170, 562]. Reconfiguration [352]. Recovery [493, 543]. Reduced [472]. Reduced-Introduction-Set [472]. Reducing [487]. Reduction [250]. Reduction-Based [250]. Reference [336, 583]. Refinement [466]. Reflections [315]. Relationships [78]. Reliable [314]. Relief [321]. Remarks [140]. Remote [463]. report [439, 507, 570, 572]. Representation [79, 344, 401, 440, 490, 547, 548]. Representations [402]. Representing [80, 82]. Requirements [172]. Research [103, 129, 144, 155, 206, 231, 254, 286, 334, 470, 486, 513–516, 528, 546, 555, 556]. Residue [201]. Resistive [452]. retrospective [558]. Reuse [543]. Reusing [355]. Reviews [109, 135, 163, 192, 215, 236, 268, 296, 420, 429].

RISC [65, 414]. Rivers [554]. Robot [69, 401, 403, 406, 510, 533–535]. Robotic [514, 540]. Robotics [508, 509, 512–516]. Robots [143]. Robustness [511]. Rochester [47]. Role [348, 551]. Routing [223]. Rover [532]. RTR [276]. Rule [359]. Rule-Based [359].

Safety [120]. Satellite [199]. Saxpy [392]. Scale [15, 49]. Schedule [359]. Scheduling [542]. Scheme [32]. Scheme-79 [32]. Science [169, 172, 439, 507]. Sciences [550]. Scientific [15, 35, 247, 438, 498, 546, 554, 555]. SDI [364]. Seamless [444]. Search [335]. Second [20, 56]. Secure [517]. Segue [492]. Self [7, 454]. Self-Checking [7]. Self-Organizing [454]. Semantic [77, 468]. Semiconductor [278, 279]. Semiconductors [551]. sequences [580]. Sequential [520, 579]. Sequoia [447]. Serious [481]. Server [374]. Set [92, 93, 472, 521]. Sets [342, 415, 547, 550]. Shapes [322]. Shared [487]. Shared-Memory [487]. Shop [310]. Shuffle [388]. Shuffle-Exchange [388]. Shuffles [44]. Sigcomm [185]. Signal [51, 398, 448]. silver [383]. Simulated [250]. Simulation [8, 321, 341, 551]. Single [330, 521]. Single-Stuck-Fault [521]. Sixth [189]. Size [375]. SLAPP [393]. Smalltalk [340, 377]. Snowbird [507]. Society [86, 98, 113, 123, 169, 219, 240, 272, 300, 409, 419, 431]. Software [61, 83, 114, 116, 120, 122, 145, 176, 177, 222, 224, 242, 251, 339, 355, 383, 413, 435, 464, 467, 494, 501, 523, 524, 526, 562]. Solid [323]. Solve [94]. Some [394, 416]. Source [449]. Soviet [478]. Space [33]. Spatial [578]. Special [1, 241]. Special-Purpose [1]. Specification [116, 578]. spectrum [571]. Speed [337, 448, 478, 559]. Spherical [402]. Spiral [464]. Spring [90, 96]. Sprite [446]. Stack [382]. Staff [139, 167, 196]. Standard [2, 16, 23–26, 527]. Standardization [347]. Standards [225, 348, 417]. Star [558].

Static [484]. Status [347]. Steering [397]. Step [396]. Stepwise [466]. Storage [4, 59–61, 336]. Storming [526]. Strategic [480, 501]. Strategies [88, 223, 352, 517]. Strategy [404]. Structure [80]. Structured [455, 466, 567]. Structures [42, 51, 323]. Structuring [304]. Stuck [521]. Study [124, 432]. Style [560]. Styles [441]. summary [572]. Super [17, 382]. Supercomputer [477, 495, 497, 572]. Supercomputers [18, 20, 148, 494]. Supercomputing [309, 554]. Superproblems [148]. Supersystem [399]. Support [118, 312, 339, 492, 522]. Surfaces [69]. Surrogate [380]. Survey [46, 48, 338, 348, 412, 414, 470, 476, 541, 545, 570]. Suspension [536]. Switching [43]. Symbolic [367, 575]. Symposium [189]. Synchronization [8, 445, 493, 545]. Synchronizing [353]. Synchronous [568]. Synergism [386]. System [10, 16, 35, 57, 170, 198, 242, 276, 303, 306, 321, 337, 349, 350, 357, 359–363, 384, 398, 402, 442, 446, 465, 467, 513, 520, 529, 540, 577]. System/38 [360]. Systems [12, 17, 38–40, 42, 43, 52, 61, 76, 81, 89, 91, 94, 105, 115, 120, 130, 141, 158, 182, 183, 209, 221, 232, 251, 263, 275, 277, 280, 291, 309, 312, 317, 332, 334, 336, 346, 363, 364, 386, 441, 443, 462, 466, 479, 481, 527, 528, 538, 569, 574]. Systolic [50, 375, 390, 392–399]. Systolic/Cellular [398].

Table [85, 112, 138, 166, 195, 218, 239, 271, 299, 408]. Tables [162, 186, 213, 262, 287, 484, 572]. talk [530]. target [88]. Task [19, 304, 439, 510]. Task-Level [510]. Taulbee [476, 570]. Taxonomic [77]. Taxonomy [489]. Technique [220]. Techniques [64, 92, 93, 248, 338, 371, 395, 433]. Technology [83, 140, 142, 170, 505, 563]. Temporal [369]. terrain [490]. terrains [534]. Test [518–520, 522].

Test-Generation [520]. Testability [6, 522]. Testbed [66–68]. Texas [411]. TEXnology [557]. Text [324]. Their [348]. Third [184, 210]. TI [498]. Tightly [447]. Time [67, 78, 119, 246, 346, 448, 481, 538]. Timing [94]. Timing-Critical [94]. Titan [477]. Toast [361]. Tolerance [281, 395]. Tolerant [273–275, 277–279, 447]. Tools [10, 84, 92, 93, 384, 553]. Top [322]. Top-Down [322]. Topology [547]. Total [198]. Trade [45, 495]. Trade-Offs [45, 495]. Traditional [468]. Transaction [447]. Transformations [537]. Translation [45]. Transport [551]. Turbulent [548]. Tutor [308]. Tutorial [201, 289, 303, 457, 485]. Twenty [189]. Twenty-Sixth [189]. Two [330]. Type [78, 200].

U [70]. U-Interpreter [70]. Ubiquitous [544]. Ultrahigh [276]. Underflow [27]. Understanding [251]. Union [478]. United [394]. Universal [579]. University [254]. Unix [10, 28, 58, 276]. Unix-Based [58]. unknown [534]. Update [105, 130, 158, 183, 209, 232, 263, 291, 418, 572]. USC [516]. Use [5, 174]. User [22, 304, 343, 354, 502, 527]. User-Interface [304]. Using [8, 44, 83, 173, 330, 340, 452, 488, 535, 539].

Valued [456, 457, 459–461, 485]. variable [580]. VAX [45, 57]. VAX-11 [45]. VAX/VMS [57]. Vector [20, 469, 547, 559]. Vehicle [536]. Verification [325, 466, 470]. Very [49, 380]. Very-Large-Scale [49]. VHSIC [329]. via [580]. Videotex [21]. Views [354, 373]. Virtual [57]. Vision [69, 318, 401, 403, 540, 575, 577]. Visual [82, 341, 358, 453, 563]. Visualization [339, 546, 549, 551–556]. Visualizing [550, 564, 565]. Visually [33]. VLIW [247]. VLSI [1, 65, 84, 95, 147, 255, 277, 290, 326, 353, 451, 458, 483, 518, 525, 567]. VMS [57]. vs [443, 469].

Wang [530]. Wars [37]. Wavefront [391]. Week [289]. Were [60]. West [289]. Whistle [519]. wide [12]. wide-band [12]. Windows [370]. Workbench [519]. Workshop [96, 184, 210, 254, 314]. Workstation [341]. world [82]. Write [465]. Write-Once [465]. Writing [345].

X [249, 497, 559]. X-MP [497, 559]. X-MP-2 [249]. Xerox [558].

Zeus [327]. ZS [542]. ZS-1 [542].

References

Foster:1980:DSV

[1] M. J. Foster and H. T. Kung. The design of special-purpose VLSI chips. Computer, 13(1):26–40, January 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coonen:1980:IGP

[2] Jerome T. Coonen. An implementation guide to a proposed standard for floating-point arithmetic. Computer, 13 (1):68–79, January 1980. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See errata in [24]. See [?, ?].

Thornton:1980:BNA

[3] J. E. Thornton. Back-end network approaches. Computer, 13(2):10–17, February 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chlamatac:1980:PIB

[4] I. Chlamatac, W. R. Franta, P. C. Patton, B., and Wells. Performance issues in back-end storage networks. Computer, 13(2):18–31, February 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Butland:1980:EGD

[5] J. Butland and S. D. Butland. An easyto-use graph drawing package. Computer, 13(2):69–80, February 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hayes:1980:TCM

[6] J. P. Hayes and E. J. McCluskey. Testability considerations in microprocessorbased design. Computer, 13(3):17– 26, March 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pradhan:1980:ECS

[7] D. K. Pradhan and J. J. Stiffler. Errorcorrecting codes and self-checking circuits. Computer, 13(3):27–37, March 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peacock:1980:DSU

[8] K. Peacock, E. Manning, and J. Wong. Synchronization of distributed simulation using broadcast algorithms. Computer, 13(4):3–10, April 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buzen:1980:MCQ

[9] Jeffrey P. Buzen and Peter J. Denning. Measuring and calculating queue length distributions. Computer, 13(4): 33–44, April 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1980:LDT

[10] Steven C. Johnson. Language development tools on the Unix system. Computer, 13(8):16–24, August 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leverett:1980:OPC

[11] Bruce W. Leverett, R. G. G. Cattell, S. O. Hobbs, J. M. Newcomer, A. H. Reiner, B. R. Schatz, and W. A. Wulf. An overview of the Production Quality Compiler-Compiler Project. Computer, 13(8):38–49, August 1980. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilkes:1980:IWL

[12] M. V. Wilkes. The impact of wideband local area communication systems on distributed computing. Computer, 13(9):22–25, September 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Isloor:1980:DPO

[13] S. S. Isloor and T. A. Marsland. The deadlock problem: An overview. Computer, 13(9):58–78, September 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flynn:1980:DIA

[14] M. Flynn. Directions and issues in architecture and language. Computer, 13(10): 5–22, October 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rodrigue:1980:PLS

[15] G. Rodrigue, E. D. Giroux, and M. Pratt. Perspectives on large-scale scientific computation. Computer, 13 (10):65–80, October 1980. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boberg:1980:PMS

[16] R. W. Boberg et al. Proposed Microcomputer System 796 Bus Standard. Computer, 13(10):89–105, October 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vick:1980:AAS

[17] C. R. Vick, S. P. Kartashev, and S. I. Kartashev. Adaptable architectures for super systems. Computer, 13(11):17– 35, November 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dennis:1980:DFS

[18] Jack B. Dennis. Data flow supercomputers. Computer, 13(11):48–56, November 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:1980:TAD

[19] W. W. Chu, L. J. Holloway, M. Lan, and K. Efe. Task allocation in distributed data processing. Computer, 13(11):57– 69, November 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kozdrowicki:1980:SGV

[20] E. W. Kozdrowicki and Douglas J. Theis. Second generation of vector supercomputers. Computer, 13(11):71– 83, November 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ball:1980:VN

[21] A. J. S. Ball, G. V. Bochmann, and J. Gecsei. Videotex networks. Computer, 13(12):8–14, December 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1981:CGR

[22] W. Myers. Computer graphics: Reaching the user. Computer, 14(3):7–10, 14, 17, March 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stevenson:1981:PSB

[23] David Stevenson. A proposed standard for binary floating-point arithmetic. Computer, 14(3):51–62, March 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coonen:1981:EIG

[24] Jerome T. Coonen. Errata: An Implementation Guide to a Proposed Standard for Floating Point Arithmetic. Computer, 14(3):62, March 1981. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See [2, ?, ?].

Cody:1981:APF

[25] William J. Cody, Jr. Analysis of proposals for the floating-point standard. Computer, 14(3):63–68, March 1981. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hough:1981:API

[26] David Hough. Applications of the proposed IEEE-754 standard for floating point arithmetic. Computer, 14(3):70– 74, March 1981. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Coonen:1981:UDN

[27] Jerome T. Coonen. Underflow and the denormalized numbers. Computer, 14 (3):75–87, March 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kernighan:1981:UPE

[28] B. W. Kernighan and J. R. Mashey. The Unix programming environment. Computer, 14(4):12–24, April 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Teitelman:1981:IPE

[29] Warren Teitelman and Larry Masinter. The Interlisp programming environment. Computer, 14(4):25–33, April 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zloof:1981:QOA

[30] M. M. Zloof. QBE/OBE: A language for office and business automation. Computer, 14(5):13–22, May 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stenning:1981:AEA

[31] V. Stenning, R. Froggatt, R. Gilbert, and E. Thomas. The Ada environment: A perspective. Computer, 14(6): 26–36, June 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sussman:1981:SLC

[32] Gerald Jay Sussman, Jack Holloway, Guy Lewis Steele, Jr., and Alan Bell.

Scheme-79 — Lisp on a chip. Computer, 14(7):10–21, July 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Webb:1981:VIM

[33] Jon A. Webb and J. K. Aggarwal. Visually interpreting the motion of objects in space. Computer, 14(8):40–46, August 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cohler:1981:FPA

[34] E. U. Cohler and J. E. Storer. Functionally parallel architecture for array processors. Computer, 14(9):28–36, September 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maron:1981:IAP

[35] N. Maron and T. A. Brengle. Integrating an array processor into a scientific computing system. Computer, 14(9):41– 44, September 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brode:1981:PFP

[36] B. Brode. Precompilation of Fortran programs to facilitate array processing. Computer, 14(9):46–51, September 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cohen:1981:HWP

[37] Danny Cohen. On Holy Wars and a plea for peace. Computer, 14(10):48– 54, October 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). This is an entertaining account of the Big-Endian and Little-Endian problems of bit and byte ordering.

Pohm:1981:CMS

[38] A. V. Pohm and T. A. Smay. Computer memory systems. Computer, 14(10):93– 110, October 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:1981:PDB

[39] S. K. Chang and T. L. Kunii. Pictorial data base systems. Computer, 14(11): 13–21, November 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:1981:PQL

[40] N. S. Chang and K. S. Fu. Picture query languages for pictorial database systems. Computer, 14(11):23– 33, November 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zobrist:1981:PIP

[41] A. L. Zobrist and G. Nagy. Pictorial information processing of LANDSAT data for geographic analysis. Computer, 14 (11):34–41, November 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chock:1981:MDS

[42] Margaret Chock, Alfonso F. Cardenas, and Allen Klinger. Manipulating data structures in pictorial information systems. Computer, 14(11):43– 50, November 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dias:1981:PSI

[43] D. M. Dias and J. R. Jump. Packet switching interconnection networks for modular systems. Computer, 14(12):43– 53, December 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:1981:INU

[44] P.-Y. Chen, D. H. Lawrie, P.-C. Yew, and D. A. Padua. Interconnection networks using shuffles. Computer, 14(12): 55–63, December 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Satyanarayanan:1981:DTV

[45] M. Satyanarayanan and D. P. Bhandarkar. Design trade-offs in VAX-11 translation buffer organization. Computer, 14(12):103–111, December 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Plattner:1981:MPE

[46] B. Plattner and J. Nievergelt. Monitoring program execution: A survey. Computer, 16(11):76–93, November 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lantz:1982:RIG

[47] K. A. Lantz, K. D. Gradischnig, J. A. Feldman, R. F., and Rashid. Rochester’s intelligent gateway. Computer, 11(10): 54–68, October 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haynes:1982:SHP

[48] L. S. Haynes, R. L. Lau, D. P. Siewiorek, and D. W. Mizel. A survey of highly parallel computing. Computer, 15(1):9– 24, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gottlieb:1982:NAV

[49] Allan Gottlieb and J. T. Schwartz. Networks and algorithms for very-large-scale parallel computation. Computer, 15(1):27–36, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kung:1982:WSA

[50] H. T. Kung. Why systolic architectures? Computer, 15(1):37–46, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmed:1982:HCC

[51] H. M. Ahmed, J. M. Delosme, and M. Morf. Highly concurrent computing structures for matrix arithmetic and signal processing. Computer, 15(1):65– 82, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Agerwala:1982:DS

[52] T. Agerwala and Arvind. DataFlow systems. Computer, 15(2):10–30, February 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ackerman:1982:DFL

[53] W. B. Ackerman. Data flow languages. Computer, 15(2):15–25, February 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davis:1982:DFP

[54] Alan L. Davis and Robert M. Keller. Data flow program graphs. Computer, 15(2):26–41, February 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Watson:1982:PDF

[55] I. Watson and J. Gurd. A practical data flow computer. Computer, 15(2): 51–57, February 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gajski:1982:SOD

[56] D. D. Gajski, D. A. Padua, D. J. Kuck, and R. H. Kuhn. A second opinion on data flow machines and languages. Computer, 15(2):58–69, February 1982. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Levy:1982:VMM

[57] H. M. Levy and P. H. Lipman. Virtual memory management in the VAX/ VMS operating system. Computer, 15 (3):35–41, March 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hwang:1982:ULC

[58] K. Hwang, W. J. Croft, G. H. Goble, B. W. Wah, F. A. Briggs, W. R. Simmons, and C. L. Coates. A Unix-Based local computer network with load balancing. Computer, 15(4):55–66, April 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kenville:1982:ODD

[59] Richard F. Kenville. Optical disk data storage. Computer, 15(7):21–25, July 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Copeland:1982:WIM

[60] George Copeland. What if mass storage were free? Computer, 15(7):27–35, July 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

OLear:1982:SCM

[61] B. T. O’Lear and J. H. Choy. Software considerations in mass storage systems. Computer, 15(7):36–44, July 1982. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawrie:1982:EAF

[62] D. H. Lawrie, J. M. Randal, and R. R. Barton. Experiments with automatic file migration. Computer, 15(7):45–55, July 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shoch:1982:EEL

[63] John F. Shoch, Y. K. Dalal, David D. Redell, and Ronald C. Crane. Evolution of the Ethernet local computer network. Computer, 15(8):10–27, August 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boral:1982:ADF

[64] H. Boral and D. J. DeWitt. Applying data flow techniques to data base machines. Computer, 15(8):57– 63, August 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Reprinted in A. R. Hurson, L. L. Miller, and S. H. Pakzad, Parallel Architectures for Database Systems, IEEE Computer Society Press, Washington, DC, 1989.

Patterson:1982:VR

[65] D. A. Patterson and C. H. Sequin. A VLSI RISC. Computer, 15(9):8–21, September 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brayer:1982:TAD

[66] K. Brayer and V. Lafleur. A testbed approach to the design of a computer communication network. Computer, 15 (10):14–23, October 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McDonald:1982:FDT

[67] W. C. McDonald and R. W. Smith. A flexible distributed testbed for realtime applications. Computer, 15(10):25– 39, October 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gehringer:1982:CT

[68] E. F. Gehringer, A. K. Jones, and Z. Z. Segall. The Cm* testbed. Computer, 15 (10):40–49, October 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hall:1982:MCS

[69] Ernest L. Hall, James K. B. Tio, Charles A. McPherson, and Firooz A. Sadjadi. Measuring curved surfaces for robot vision. Computer, 15(12):42–54, December 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arvind:1982:U

[70] Arvind and Kim P. Gostelow. The U-Interpreter. Computer, 16(1):42– 49, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fu:1983:ICA

[71] K. Hwang and K. Fu. Integrated computer architectures for image processing and database management. Computer, 16(1):51–60, January 1983. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atkinson:1983:APP

[72] M. P. Atkinson, P. J. Bailey, K. J. Chisholm, P. W. Cockshott, and R. Morrison. An approach to persistent programming. Computer, 16(4):360– 365, April 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Benhamou:1983:MIG

[73] Eric Benhamou and Judy Estrin. Multilevel internetworking gateways: Architecture and applications. Computer, 16 (9):27–34, September 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hinden:1983:DII

[74] R. Hinden, J. Haverty, and A. Sheltzer. The DARPA Internet: Interconnecting heterogeneous computer networks with gateways. Computer, 16(9):38–48, September 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). papers box.

Thornton:1983:HNL

[75] James E. Thornton and Gary S. Christensen. Hyperchannel network links. Computer, 16(9):50–54, September 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Redell:1983:IEM

[76] David D. Redell and James E. White. Interconnecting electronic mail systems. Computer, 16(9):55–63, September 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brachman:1983:WII

[77] Ronald J. Brachman. What IS-A is and isn’t: An analysis of taxonomic links in semantic networks. Computer, 16 (10):30–36, October 1983. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schubert:1983:DTP

[78] Lenhart K. Schubert, Mary Angela Papalaskaris, and Jay Taugher. Determining type, part, color, and time relationships. Computer, 16(10):53– 60, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Special Issue on Knowledge Representation.

Brachman:1983:KAF

[79] Ronald J. Brachman, Richard E. Fikes, and Hector J. Levesque. KRYPTON: A functional approach to knowledge representation. Computer, 16(10):67– 73, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Special Issue on Knowledge Representation.

Davis:1983:RSB

[80] Randall Davis and Howard Shrobe. Representing structure and behavior of digital hardware. Computer, 16(10):75– 81, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mylopoulos:1983:BKS

[81] John Mylopoulos, Tetsutaro Shibahara, and John K. Tsotsos. Building knowledge-based systems: The PSN experience. Computer, 16(10):83–89, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Havens:1983:RKV

[82] W. Havens and A. Mackworth. Representing knowledge of the visual world. Computer, 16(10):90–96, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balzer:1983:STU

[83] Robert Balzer, Thomas E. Cheatham, Jr., and Cordell Green. Software technology in the 1990s: Using a new paradigm. Computer, 16(11):39–45, November 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gajski:1983:GEI

[84] D. Gajski and R. Kuhn. Guest Editors’ introduction: New VLSI tools. Computer, 16(12):11–14, December 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCa

[85] Anonymous. Table of contents. Computer, 17(1):1, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICSa

[86] Anonymous. IEEE Computer Society. Computer, 17(1):3, January 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:LEa

[87] Anonymous. Letters to the editor. Computer, 17(1):5, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1984:BPS

[88] J. K. F. Lee and A. J. Smith. Branch prediction strategies and branch target buffer design. Computer, 17(1):6– 22, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wah:1984:FPD

[89] ?. Wah. File placement on distributed computer systems. Computer, 17(1):23– 32, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CS

[90] Anonymous. Compcon Spring 84. Computer, 17(1):32a–32h, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gligor:1984:IHD

[91] V. D. Gligor and G. L. Luckenbaugh. Interconnecting heterogeneous database management systems. Computer, 17 (1):33–43, January 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boari:1984:MMP

[92] M. Boari, S. Crespi-Reghizzi, A. Dapra, and A. Natali. Multiple-microprocessor programming techniques: MML, a new set of tools. Computer, 17(1):47– 59, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boari:1984:MPT

[93] Maurelio Boari, Stefano Crespi-Reghizzi, Alberto Dapra, Francesco Maderna, and Antonio Natali. Multiplemicroprocessor programming techniques: MML, a new set of tools. Computer, 17(1):47–59, January 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ma:1984:MST

[94] ?. Ma. A model to solve timing-critical application problems in distributed computer systems. Computer, 17(1):62– 68, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grinberg:1984:CVA

[95] J. Grinberg, G. R. Nudd, and R. D. Etchells. A cellular VLSI architecture. Computer, 17(1):69–81, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balde:1984:ICPb

[96] ?. Balde. IEEE Computer Packaging Committee spring packaging workshop. Computer, 17(1):83–86, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matley:1984:OC

[97] B. G. Matley and J. A. Grosberg. The open channel. Computer, 17(1):87– 88, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICSb

[98] Anonymous. IEEE Computer Society. Computer, 17(1):88a–88p, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1984:NPa

[99] D. A. Michalopoulos. New products. Computer, 17(1):89–97, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NLa

[100] Anonymous. New literature. Computer, 17(1):97, January 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RIAa

[101] Anonymous. Recent IC announcements. Computer, 17(1):100, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RMAa

[102] Anonymous. Recent microsystems announcements. Computer, 17(1):103, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NARa

[103] Anonymous. New applications & recent research. Computer, 17(1):104– 105, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Ea

[104] Anonymous. Erratum. Computer, 17(1):

106, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:UICa

[105] Anonymous. Update: information for the computer systems design professional. Computer, 17(1):106–113, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Cb

[106] Anonymous. Calendar. Computer, 17 (1):114–118, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPa

[107] Anonymous. Call for papers. Computer, 17(1):115–116, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CAa

[108] Anonymous. Classified ads. Computer, 17(1):119–131, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:BRa

[109] Anonymous. Book reviews. Computer, 17(1):132–133, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Ba

[110] Anonymous. The bookshelf. Computer, 17(1):135, January 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FCa

[111] Anonymous. [Front cover]. Computer, 17(1):c1–c2, January 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCb

[112] Anonymous. Table of contents. Computer, 17(2):1, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICSc

[113] Anonymous. IEEE Computer Society.

Computer, 17(2):4, February 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lauber:1984:GEI

[114] R. J. Lauber. Guest Editor’s introduction: Software for industrial process control. Computer, 17(2):6–8, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schoeffler:1984:DCS

[115] ?. Schoeffler. Distributed computer systems for industrial process control. Computer, 17(2):11–18, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walter:1984:CSS

[116] ?. Walter. Control software specification and design: an overview. Computer, 17 (2):20–23, February 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CFa

[117] Anonymous. Compcon Fall 84. Computer, 17(2):24–25, February 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weide:1984:PCI

[118] B. W. Weide, M. E. Brown, J. Ramanathan, and K. Schwan. Process control: Integration and design methodology support. Computer, 17(2):27– 32, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Steusloff:1984:ART

[119] ?. Steusloff. Advanced real-time languages for distributed industrial process control. Computer, 17(2):37–46, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leveson:1984:SSC

[120] ?. Leveson. Software safety in computercontrolled systems. Computer, 17(2):48– 55, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FICa

[121] Anonymous. The First International Conference on Computers and Applications. Computer, 17(2):56–57, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matsumoto:1984:MIS

[122] ?. Matsumoto. Management of industrial software production. Computer, 17 (2):59–72, February 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICSd

[123] Anonymous. IEEE Computer Society 1984 Conference on Ada Applications and Environments. Computer, 17(2):

74, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bentley:1984:CSA

[124] ?. Bentley. A case study in applied algorithm design. Computer, 17(2):75– 88, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1984:NPb

[125] D. A. Michalopoulos. New products.

Computer, 17(2):90–95, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NLb

[126] Anonymous. New literature. Computer, 17(2):95, February 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RIAb

[127] Anonymous. Recent IC announcements.

Computer, 17(2):96, February 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RMAb

[128] Anonymous. Recent microsystem announcements. Computer, 17(2):98, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NARb

[129] Anonymous. New applications & recent research. Computer, 17(2):102– 103, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:UICb

[130] Anonymous. Update: information for the computer systems design professional. Computer, 17(2):104–110, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CAb

[131] Anonymous. Classified ads. Computer, 17(2):111–119, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Cc

[132] Anonymous. Calendar. Computer, 17 (2):120–124, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPb

[133] Anonymous. Call for papers. Computer, 17(2):121–122, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICDa

[134] Anonymous. International conference on data engineering. Computer, 17(2):125– 130, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:BRb

[135] Anonymous. Book reviews. Computer, 17(2):131–134, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Bb

[136] Anonymous. The bookshelf. Computer, 17(2):135, February 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FCb

[137] Anonymous. [Front cover]. Computer, 17(2):c1–c2, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCc

[138] Anonymous. Table of contents. Computer, 17(3):1, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICMa

[139] Anonymous. [IEEE Computer — magazine’s staff]. Computer, 17(3):2, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yau:1984:ECR

[140] ?. Yau. Editor-in-Chief’s remarks:

Japanese computer technology and culture. Computer, 17(3):4–5, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moto-oka:1984:FGC

[141] ?. Moto-oka and ?. Stone. Fifthgeneration computer systems: a Japanese project. Computer, 17(3):6– 13, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuchi:1984:JAH

[142] ?. Fuchi, ?. Sato, and ?. Miller. Japanese approaches to high-technology R&D. Computer, 17(3):14–18, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Togai:1984:JNG

[143] ?. Togai. Japan’s next generation of robots. Computer, 17(3):19–25, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stone:1984:CRJ

[144] ?. Stone. Computer research in Japan. Computer, 17(3):26–32, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tajima:1984:IJS

[145] ?. Tajima and ?. Matsubara. Inside the Japanese software industry. Computer, 17(3):34–43, March 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Treleaven:1984:FCL

[146] ?. Treleaven and ?. Lima. Future computers: Logic, data flow, …, control flow? Computer, 17(3):47–58, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:IICa

[147] Anonymous. IEEE International Conference on Computer Design: VLSI in Computers. Computer, 17(3):61, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Norrie:1984:SSA

[148] ?. Norrie. Supercomputers for superproblems: an architectural introduction. Computer, 17(3):62–74, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:IICb

[149] Anonymous. IEEE Infocom 85: Computers and communications integration: The confluence at mid-decade. Computer, 17(3):72, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hartley:1984:CCF

[150] ?. Hartley. CRIB: Computer faultfinding through knowledge engineering. Computer, 17(3):76–83, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sen:1984:OC

[151] Chandan Sen and R. G. Estell. The open channel. Computer, 17(3):87, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carlson:1984:IC

[152] ?. Carlson and ?. Simmons. Intersociety cooperation. Computer, 17(3):88– 89, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1984:NPd

[153] D. A. Michalopoulos. New products.

Computer, 17(3):90–97, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NLc

[154] Anonymous. New literature. Computer, 17(3):97, March 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NARc

[155] Anonymous. New applications & recent research. Computer, 17(3):98–99, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RIAc

[156] Anonymous. Recent IC announcements. Computer, 17(3):100, March 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RMAc

[157] Anonymous. Recent microsystem announcements. Computer, 17(3):101, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:UICc

[158] Anonymous. Update: information for the computer systems design professional. Computer, 17(3):102–109, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CAc

[159] Anonymous. Classified ads. Computer, 17(3):110–115, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Cd

[160] Anonymous. Calendar. Computer, 17 (3):116–122, March 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPc

[161] Anonymous. Call for papers. Computer, 17(3):117–118, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCd

[162] Anonymous. Tables of contents. Computer, 17(3):123, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:BRc

[163] Anonymous. Book reviews. Computer, 17(3):124–126, March 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Bc

[164] Anonymous. The bookshelf. Computer, 17(3):127, March 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FCc

[165] Anonymous. [Front cover]. Computer, 17(3):c1–c2, March 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCe

[166] Anonymous. Table of contents. Computer, 17(4):1, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICMb

[167] Anonymous. [IEEE Computer — magazine’s staff]. Computer, 17(4):2, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Booth:1984:GEI

[168] T. L. Booth and J. T. Cain. Guest Editor’s introduction: Computers in education: a new learning environment. Computer, 17(4):4–6, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cain:1984:ICS

[169] ?. Cain, ?. Langdon, and ?. Varanasi. The IEEE Computer Society model program in computer science and engineering. Computer, 17(4):8–17, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RGA

[170] Anonymous. Recommendations and guidelines for associate degree programs in computer system technology. Computer, 17(4):18–22, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1984:ACP

[171] ?. Jones and ?. Mulder. Accreditation in the computer profession. Computer, 17(4):24–27, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mulder:1984:CSP

[172] ?. Mulder and ?. Dalphin. Computer science program requirements and accreditation. Computer, 17(4):30–35, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bray:1984:UPC

[173] ?. Bray. Using personal computers at the college level. Computer, 17(4):36– 43, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rogers:1984:CUP

[174] ?. Rogers. Computer use in precollege education. Computer, 17(4):46–52, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moore:1984:GAL

[175] ?. Moore, ?. Geer, and ?. Graf. Gridnet:

an alternative large distributed network. Computer, 17(4):57–66, April 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Albert:1984:CSP

[176] ?. Albert and ?. Morse. Combatting software piracy by encryption and key management. Computer, 17(4):68–73, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bryan:1984:PAI

[177] ?. Bryan and ?. Siegel. Product assurance: Insurance against a software disaster. Computer, 17(4):75–83, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1984:NPc

[178] D. A. Michalopoulos. New products.

Computer, 17(4):84–93, April 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NLd

[179] Anonymous. New literature. Computer, 17(4):93, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RIAd

[180] Anonymous. Recent IC announcements.

Computer, 17(4):95, April 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RMAd

[181] Anonymous. Recent microsystem announcements. Computer, 17(4):96, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICDb

[182] Anonymous. The 4th International Conference on Distributed Computing Systems. Computer, 17(4):96a–96h, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:UICd

[183] Anonymous. Update: information for the computer systems design professional. Computer, 17(4):97–104, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TAWa

[184] Anonymous. Third Annual Workshop on Interactive Computing: CAE Electrical Engineering Education. Computer, 17(4):105, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:AS

[185] Anonymous. ACM Sigcomm ’84. Computer, 17(4):106–107, April 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCf

[186] Anonymous. Tables of contents. Computer, 17(4):108–109, April 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Ce

[187] Anonymous. Calendar. Computer, 17 (4):110–116, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPd

[188] Anonymous. Call for papers. Computer, 17(4):111–112, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TSI

[189] Anonymous. Twenty-Sixth International Symposium Applications of Microcomputers. Computer, 17(4):117, April 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CAd

[190] Anonymous. Classified ads. Computer, 17(4):118–121, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FE

[191] Anonymous. Future events. Computer, 17(4):122–123, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:BRd

[192] Anonymous. Book reviews. Computer, 17(4):124–126, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Bd

[193] Anonymous. The bookshelf. Computer, 17(4):127, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FCd

[194] Anonymous. [Front cover]. Computer, 17(4):c1–c2, April 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCg

[195] Anonymous. Table of contents. Computer, 17(5):1, May 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:LS

[196] Anonymous. List of staff. Computer, 17(5):2, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rich:1984:GEA

[197] ?. Rich. The gradual expansion of artificial intelligence. Computer, 17(5): 4–12, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roman:1984:TSD

[198] ?. Roman, ?. Stucki, ?. Ball, and ?. Gillett. A total system design framework. Computer, 17(5):15–26, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Franta:1984:HLN

[199] ?. Franta and ?. Heath. Hyperchannel local network interconnection through satellite links. Computer, 17(5):30– 39, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flowers:1984:DTM

[200] Jim Flowers. Digital type manufacture: an interactive approach. Computer, 17(5):40–48, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Taylor:1984:RAT

[201] Fred J. Taylor. Residue arithmetic: a tutorial with examples. Computer, 17(5): 50–62, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roussopoulos:1984:AMD

[202] N. Roussopoulos and R. T. Yeh. An adaptable methodology for database design. Computer, 17(5):64–80, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jensen:1984:OC

[203] A. S. Jensen. The open channel. Computer, 17(5):81–82, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1984:NPe

[204] D. A. Michalopoulos. New products.

Computer, 17(5):84–97, May 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NLe

[205] Anonymous. New literature. Computer, 17(5):97, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NARd

[206] Anonymous. New applications & recent research. Computer, 17(5):98–99, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RIAe

[207] Anonymous. Recent IC announcements. Computer, 17(5):100, May 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RMAe

[208] Anonymous. Recent microsystem announcements. Computer, 17(5):101, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:UICe

[209] Anonymous. Update: information for the computer systems design professional. Computer, 17(5):102–109, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TAWb

[210] Anonymous. Third Annual Workshop on Interactive Computing: CAE Electrical Engineering Education. Computer, 17(5):104, May 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Cg

[211] Anonymous. Calendar. Computer, 17 (5):110, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPe

[212] Anonymous. Call for papers. Computer, 17(5):111–116, May 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCh

[213] Anonymous. Tables of contents. Computer, 17(5):117–119, May 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CAe

[214] Anonymous. Classified ads. Computer, 17(5):121–123, May 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:BRe

[215] Anonymous. Book reviews. Computer, 17(5):125–126, May 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Be

[216] Anonymous. The bookshelf. Computer, 17(5):127, May 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FCe

[217] Anonymous. [Front cover]. Computer, 17(5):c1–c2, May 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCi

[218] Anonymous. Table of contents. Computer, 17(6):1–103, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICSe

[219] Anonymous. IEEE Computer Society.

Computer, 17(6):3–4, June 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Welch:1984:THP

[220] Terry A. Welch. A technique for highperformance data compression. Computer, 17(6):8–19, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shatz:1984:CMP

[221] ?. Shatz. Communication mechanisms for programming distributed systems. Computer, 17(6):21–28, June 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boehm:1984:SDE

[222] ?. Boehm, ?. Penedo, ?. Stuckle, ?. Williams, and ?. Pyster. A software development environment for improving productivity. Computer, 17(6): 30–44, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hsieh:1984:RSC

[223] W. N. Hsieh and I. Gitman. Routing strategies in computer networks. Computer, 17(6):46–56, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zelkowitz:1984:SEP

[224] ?. Zelkowitz, ?. Yeh, ?. Hamlet, ?. Gannon, and ?. Basili. Software engineering practices in the US and Japan. Computer, 17(6):57–66, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bell:1984:SCH

[225] ?. Bell. Standards can help us. Computer, 17(6):71–79, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

James:1984:OC

[226] G. James. The open channel. Computer, 17(6):83–84, June 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1984:NPf

[227] D. A. Michalopoulos. New products.

Computer, 17(6):86–91, June 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NLf

[228] Anonymous. New literature. Computer, 17(6):91, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RIAf

[229] Anonymous. Recent IC announcements.

Computer, 17(6):92, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RMAf

[230] Anonymous. Recent microsystem announcements. Computer, 17(6):93, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NARe

[231] Anonymous. New applications & recent research. Computer, 17(6):94, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:UICf

[232] Anonymous. Update: information for the computer systems design professional. Computer, 17(6):96–101, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CAf

[233] Anonymous. Classified ads. Computer, 17(6):104–106, June 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Ch

[234] Anonymous. Calendar. Computer, 17 (6):108–115, June 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPf

[235] Anonymous. Call for papers. Computer, 17(6):109–110, June 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:BRf

[236] Anonymous. Book reviews. Computer, 17(6):116–117, June 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Bf

[237] Anonymous. The bookshelf. Computer, 17(6):118–119, June 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FCf

[238] Anonymous. [Front cover]. Computer, 17(6):c1–c2, June 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCj

[239] Anonymous. Table of contents. Computer, 17(7):3, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICSf

[240] Anonymous. IEEE Computer Society.

Computer, 17(7):4, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yau:1984:SM

[241] S. S. Yau. Special message. Computer, 17(7):8–11, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Martin:1984:GEI

[242] ?. Martin and ?. Brice. Guest Editors’ introduction: Effect of hardwaresoftware interaction on system performance. Computer, 17(7):12–13, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

MacDougall:1984:ILP

[243] ?. MacDougall. Instruction-level program and processor modeling. Computer, 17(7):14–24, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

MacDougall:1984:IPP

[244] Myron MacDougall. Instruction-level program and processor modeling. Computer, 17(7):14–24, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Snyder:1984:PPP

[245] L. Snyder. Parallel programming and the Poker programming environment. Computer, 17(7):27–36, July 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adams:1984:MAE

[246] ?. Adams and ?. Crockett. Modeling algorithm execution time on processor arrays. Computer, 17(7):38–43, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fisher:1984:VMM

[247] ?. Fisher. The VLIW machine: a multiprocessor for compiling scientific code. Computer, 17(7):45–53, July 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Babb:1984:PPL

[248] ?. Babb. Parallel processing with largegrain data flow techniques. Computer, 17(7):55–61, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Larson:1984:MCX

[249] ?. Larson. Multitasking on the Cray X-MP-2 multiprocessor. Computer, 17(7): 62–69, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keller:1984:SPR

[250] R. M. Keller and F. C. H. Lin. Simulated performance of a reduction-based multiprocessor. Computer, 17(7):70–82, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Browne:1984:UEB

[251] ?. Browne. Understanding execution behavior of software systems. Computer, 17(7):83–87, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fernbach:1984:OC

[252] S. Fernbach, R. B. Hibbs, and J. Horning. The open channel. Computer, 17(7): 89–90, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPg

[253] Anonymous. Call for papers. Computer, 17(7):91, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Browne:1984:UIG

[254] J. C. Browne. University, industry, and government collaboration workshop: Research in parallel computing. Computer, 17(7):92–93, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:IICc

[255] Anonymous. IEEE international conference on computer design: VLSI in computers. Computer, 17(7):95–106, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NPa

[256] Anonymous. New products. Computer, 17(7):107–115, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICP

[257] Anonymous. 1984 international conference on parallel processing. Computer, 17(7):112–113, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NLg

[258] Anonymous. New literature. Computer, 17(7):115, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RIAg

[259] Anonymous. Recent IC announcements. Computer, 17(7):116, July 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RMAg

[260] Anonymous. Recent microsystem announcements. Computer, 17(7):117, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NA

[261] Anonymous. New applications. Computer, 17(7):118–119, July 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCk

[262] Anonymous. Tables of contents. Computer, 17(7):120–121, July 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:UICg

[263] Anonymous. Update: information for the computer systems design professional. Computer, 17(7):123–131, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Eb

[264] Anonymous. Erratum. Computer, 17 (7):125, July 1984. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Ci

[265] Anonymous. Calendar. Computer, 17 (7):132–137, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPh

[266] Anonymous. Call for papers. Computer, 17(7):133–134, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CAg

[267] Anonymous. Classified ads. Computer, 17(7):138–140, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:BRg

[268] Anonymous. Book reviews. Computer, 17(7):141–142, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Bg

[269] Anonymous. The bookshelf. Computer, 17(7):143, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FCg

[270] Anonymous. [Front cover]. Computer, 17(7):c1–c2, July 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCl

[271] Anonymous. Table of contents. Computer, 17(8):3, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICSg

[272] Anonymous. IEEE Computer Society. Computer, 17(8):4, August 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Toy:1984:GEI

[273] ?. Toy and ?. Morganti. Guest Editors’ introduction: Fault-tolerant computing. Computer, 17(8):6–7, August 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siewiorek:1984:AFT

[274] ?. Siewiorek. Architecture of faulttolerant computers. Computer, 17(8): 9–18, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serlin:1984:FTS

[275] ?. Serlin. Fault-tolerant systems in commercial applications. Computer, 17(8): 19–30, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wallace:1984:DUA

[276] ?. Wallace and ?. Barnes. Designing for ultrahigh availability: The Unix RTR operating system. Computer, 17(8):31– 39, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1984:IVA

[277] ?. Johnson. The Intel 432: a VLSI architecture for fault-tolerant computer systems. Computer, 17(8):40–48, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sarrazin:1984:FSM

[278] D. B. Sarrazin and M. Malek. Faulttolerant semiconductor memories. Computer, 17(8):49–56, August 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sarrazin:1984:FTS

[279] ?. Sarrazin and ?. Malek. Fault-tolerant semiconductor memories. Computer, 17 (8):49–56, August 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ihara:1984:ADC

[280] ?. Ihara and ?. Mori. Autonomous decentralized computer control systems. Computer, 17(8):57–66, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Avizienis:1984:FTD

[281] ?. Avizienis and ?. Kelly. Fault tolerance by design diversity: Concepts and experiments. Computer, 17(8):67–80, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NPb

[282] Anonymous. New products. Computer, 17(8):81–89, August 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NLh

[283] Anonymous. New literature. Computer, 17(8):89, August 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RIAh

[284] Anonymous. Recent IC announcements. Computer, 17(8):90, August 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:RMAh

[285] Anonymous. Recent microsystem announcements. Computer, 17(8):91, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:NARf

[286] Anonymous. New applications & recent research. Computer, 17(8):94– 95, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCm

[287] Anonymous. Tables of contents. Computer, 17(8):96, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CFb

[288] Anonymous. Compcon 84 Fall. Computer, 17(8):97–104, August 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TWW

[289] Anonymous. Tutorial Week West 84. Computer, 17(8):105–112, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:IICd

[290] Anonymous. IEEE international conference on computer design: VLSI in computers. Computer, 17(8):113– 124, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:UICh

[291] Anonymous. Update: information for the computer systems design professional. Computer, 17(8):125–132, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Ec

[292] Anonymous. Erratum. Computer, 17(8):

128, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CAh

[293] Anonymous. Classified ads. Computer, 17(8):133–135, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Cj

[294] Anonymous. Calendar. Computer, 17 (8):136–140, August 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPi

[295] Anonymous. Call for papers. Computer, 17(8):137–138, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:BRh

[296] Anonymous. Book reviews. Computer, 17(8):141–142, August 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Bh

[297] Anonymous. The bookshelf. Computer, 17(8):143, August 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FCh

[298] Anonymous. [Front cover]. Computer, 17(8):c1–c2, August 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:TCn

[299] Anonymous. Table of contents. Computer, 17(9):3, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:ICSh

[300] Anonymous. IEEE Computer Society. Computer, 17(9):4–5, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:LEb

[301] Anonymous. Letters to the editor. Computer, 17(9):6–7, September 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karna:1984:GEI

[302] K. N. Karna. Guest Editor’s introduction: Artificial intelligence for manmachine interface. Computer, 17(9):8– 9, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hayes-Roth:1984:KBEa

[303] ?. Hayes-Roth. The knowledge-based expert system: a tutorial. Computer, 17(9):11–28, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1984:DTD

[304] ?. Smith, ?. Lafue, ?. Schoen, and ?. Vestal. Declarative task description as a user-interface structuring mechanism. Computer, 17(9):29–38, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rich:1984:NLI

[305] ?. Rich. Natural-language interfaces.

Computer, 17(9):39–47, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mittal:1984:PAK

[306] S. Mittal, B. Chandrasekaran, and J. Sticklen. Patrec: A knowledgedirected database for a diagnostic expert system. Computer, 17(9):51–58, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mittal:1984:PKD

[307] ?. Mittal, ?. Chandrasekaran, and ?. Sticklen. Patrec: a knowledgedirected database for a diagnostic expert. Computer, 17(9):51–58, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woolf:1984:BCT

[308] ?. Woolf and ?. McDonald. Building a computer tutor: Design issues. Computer, 17(9):61–73, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:FICb

[309] Anonymous. First international conference on supercomputing systems. Computer, 17(9):74–75, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bourne:1984:AMA

[310] ?. Bourne and ?. Fox. Autonomous manufacturing: Automating the jobshop. Computer, 17(9):76–86, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:CPj

[311] Anonymous. Call for papers. Computer, 17(9):87, September 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dutta:1984:AIA

[312] ?. Dutta and ?. Basu. An artificial intelligence approach to model management in decision support systems. Computer, 17(9):89–97, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1984:Ck

[313] Anonymous. Compsac 84. Computer, 17(9):99–107, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kayton:1984:RPC

[314] M. Kayton. Reliable Process Controls Workshop. Computer, 17(9):108–110, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parrish:1984:FIC

[315] ?. Parrish. The First International Conference on Computers and Applications: Reflections. Computer, 17(9):111–113, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

MacHover:1984:ICG

[316] C. MacHover and W. Myers. Interactive computer graphics. Computer, 17(10): 145–161, October 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1984:AOS

[317] R. L. Brown and P. J. Denning. Advanced operating systems. Computer, 17(10):173–190, October 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fu:1984:PRC

[318] K. S. Fu and A. Rosenfeld. Pattern recognition and computer vision. Computer, 17(10):274–282, October 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Glinert:1984:PIG

[319] E. P. Glinert and S. L. Tanimoto. PICT:

an interactive graphical programming environment. Computer, 17(11):7–25, November 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shemer:1984:GDC

[320] J. Shemer and P. M. Neches. The genesis of a database computer. Computer, 17(11):42–??, November 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laurent:1984:GAG

[321] D. Laurent and S. Motet. GEOMATIC: A 3-D graphic relief simulation system. Computer, 17(12):25–30, December 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yoshiura:1984:TCM

[322] H. Yoshiura, K. Fujimura, and T. L. Kunii. Top-down construction of 3-D mechanical object shapes from engineering drawings. Computer, 17(12):32– 40, December 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woo:1984:ISM

[323] T. C. Woo. Interfacing solid modeling to CAD and CAM: Data structures and algorithms for decomposing a solid. Computer, 17(12):44–49, December 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:1985:CKT

[324] Y. Chu. Chinese/Kanji text and data processing. Computer, 18(1):10– 66, January 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maruyama:1985:HV

[325] F. Maruyama and M. Fujita. Hardware verification. Computer, 18(2): 22–32, February 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). U. S. Copyright Clearance Center Code: 0018-9162/85/0200-0022$01.00.

Suzuki:1985:CPE

[326] N. Suzuki. Concurrent PROLOG as an efficient VLSI design language. Computer, 18(2):33–40, February 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). U. S. Copyright Clearance Center Code: 0018-9162/85/0200-0033$01.00.

Lieberherr:1985:ZAL

[327] K. J. Lieberherr and S. M. German.

Zeus: A language for expressing algorithms in hardware. Computer, 18(2): 55–65, February 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Zeus.

Piloty:1985:CPC

[328] R. Piloty and D. Borrione. The CON-LAN project: concepts, implementations and applications. Computer, 18 (2):81–92, February 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shahdad:1985:VHD

[329] M. Shahdad, R. Lipsett, F. E.

Marschner, K. Sheehan, H. Cohen, R. Waxman, and D. Ackley. VHSIC hardware description language. Computer, 18(2):94–103, February 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). VHDL.

Berman:1985:CRT

[330] S. Berman, P. Parikh, and C. S. G. Lee.

Computer recognition of two overlapping parts using a single camera. Computer, 18(3):70–80, March 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1985:NCA

[331] Anonymous. New configuration added, price cut, for processor. Computer, 18 (5):122–??, May 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gajski:1985:EIM

[332] Daniel D. Gajski and Jih-Kwon Peir. Essential issues in multiprocessor systems. Computer, 18(6):9–27, June 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schneck:1985:PPP

[333] P. B. Schneck, D. Austin, S. L. Squires, J. Lehmann, D. Mizell, and K. Wallgren. Parallel processor programs in the Federal Government. Computer, 18(6): 43–56, June 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Murakami:1985:RPM

[334] K. Murakami, T. Kakuta, R. Onai, and N. Ito. Research on parallel machine architecture for fifth-generation computer systems. Computer, 18(6): 76–92, June 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Month June.

Wah:1985:MCS

[335] Benjamin W. Wah, G. J. Li, and C. F. Yu. Multiprocessing of combinatorial search problems. Computer, 18(6):93– 108, June 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:1985:TRM

[336] S. W. Miller and M. W. Collins. Toward a reference model of mass storage systems. Computer, 18(7):9–22, July 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ammon:1985:HLJ

[337] G. Ammon, J. Calabria, and D. Thomas.

A high-speed, large-capacity, jukebox optical disk system. Computer, 18(7): 36–46, July 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raeder:1985:SCG

[338] G. Raeder. A survey of current graphical programming techniques. Computer, 18(8):11–25, August 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1985:PVG

[339] G. P. Brown, R. T. Carling, C. F. Herot, D. A. Kramlich, and P. Souza. Program visualization: Graphical support for software development. Computer, 18 (8):27–35, August 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

London:1985:APU

[340] R. L. London and R. A. Duisberg. Animating programs using Smalltalk. Computer, 18(8):61–71, August 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Melamed:1985:VSP

[341] B. Melamed and R. J. T. Morris. Visual simulation: The performance analysis workstation. Computer, 18(8):87– 94, August 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Colwell:1985:ISB

[342] Robert P. Colwell, Charles Y. Hitchcock, III, E. Douglas Jensen, H. M. Brinkley Sprunt, and Charles P. Kollar. Instruction sets and beyond: Computers, complexity, and controversy. Computer, 18 (9):8–19, September 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coutaz:1985:AUI

[343] Joëlle Coutaz. Abstractions for user interface design. Computer, 18(9):21– 34, September 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hartzband:1985:EKR

[344] D. J. Hartzband and F. J. Maryanski. Enhancing knowledge representation in engineering databases. Computer, 18(9): 39–??, September 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yankelovich:1985:RWE

[345] Nicole Yankelovich, Norman Meyrowitz, and Andries van Dam. Reading and writing the electronic book. Computer, 18(10):15–30, October 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sarin:1985:CRC

[346] S. Sarin and I. Greif. Computer-based real-time conferencing systems. Computer, 18(10):33–45, October 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horak:1985:ODA

[347] Wolfgang Horak. Office Document Architecture and Office Document Interchange formats: Current status of international standardization. Computer, 18 (10):50–60, October 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bono:1985:SGS

[348] Peter R. Bono. A survey of graphics standards and their role in information interchange. Computer, 18(10):63– 75, October 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reynolds:1985:DEM

[349] J. K. Reynolds, J. B. Postel, A. R.

Katz, G. G. Finn, and A. L. DeSchon. The DARPA experimental multimedia mail system. Computer, 18(10):82– 91, October 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sakata:1985:DIM

[350] S. Sakata and T. Ueda. A distributed interoffice mail system. Computer, 18(10): 106–116, October 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carlson:1985:APD

[351] W. W. Carlson and K. Hwang. Algorithmic performance of dataflow multiprocessors. Computer, 18(12):30–40, December 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yalamanchili:1985:RSP

[352] S. Yalamanchili and J. K. Aggarwal. Reconfiguration strategies for parallel architectures. Computer, 18(12):44–61, December 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fisher:1985:SLV

[353] A. L. Fisher and H. T. Kung. Synchronizing large VLSI processor arrays. Computer, C-34(8):734–740, August 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Navathe:1986:IUV

[354] S. Navathe, R. Elmasri, and J. Larson. Integrating user views in database design. Computer, 19(1):50–62, January 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goguen:1986:RIS

[355] Joseph Goguen. Reusing and interconnecting software components. Computer, 19(2):16–28, February 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Reprinted in Tutorial: Software Reusability, Peter Freeman, editor, IEEE Computer Society, 1987, pages 251–263, and in Domain Analysis and Software Systems Modelling, Ruben Prieto-Diaz and Guillermo Arango, editors, IEEE Computer Society, 1991, pages 125–137.

Agrawal:1986:EPM

[356] D. P. Agrawal, V. K. Janakiram, and G. C. Pathak. Evaluating the performance of multicomputer configurations.

Computer, 19(5):23–37, May 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fisher:1986:IDD

[357] D. A. Fisher and R. M. Weatherly. Issues in the design of a distributed operating system for Ada. Computer, 19(5): 38–47, May 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Larson:1986:VAB

[358] James A. Larson. A visual approach to browsing in a database environment. Computer, 19(6):62–71, June 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bruno:1986:RSS

[359] G. Bruno, A. Elia, and P. Laface. A rulebased system to schedule production. Computer, 19(7):32–??, July 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stroebel:1986:CPE

[360] G. J. Stroebel, R. D. Baxter, and M. J.

Denney. A capacity planning expert system for IBM System/38. Computer, 19 (7):42–??, July 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Talukdar:1986:TPS

[361] S. N. Talukdar, E. Cardozo, and L. V.

Leao. Toast: The power system operator’s assistant. Computer, 19(7): 53–??, July 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1986:KCM

[362] David C. Brown and B. Chandrasekaran. Knowledge and control for a mechanical design expert system. Computer, 19(7):92–100, July 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mittal:1986:PES

[363] S. Mittal, C. L. Dym, and M. Morjaria. Pride: An expert system for the design of paper handling systems. Computer, 19(7):102–??, July 1986. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chien:1986:ESS

[364] Y.-T. Chien and J. Liebowitz. Expert systems in the SDI environment. Computer, 19(7):115–??, July 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mundie:1986:PPA

[365] D. A. Mundie and D. A. Fisher. Parallel processing in Ada. Computer, 19(8): 20–25, August 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahuja:1986:DPL

[366] Sudhir Ahuja, Nicholas Carriero, and David Gelernter. Domesticating parallelism: Linda and friends. Computer, 19 (8):26–34, August 1986. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Halstead:1986:PSC

[367] R. H. Halstead. Parallel symbolic computing. Computer, 19(8):35–43, August 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hudak:1986:PP

[368] P. R. Hudak. Para-functional programming. Computer, 19(8):60–70, August 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Snodgrass:1986:TD

[369] R. T. Snodgrass and I. Ahn. Temporal databases. Computer, 19(9):35–42, September 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1986:CEI

[370] B. A. Myers. A complete and efficient implementation of covered windows. Computer, 19(9):57–67, September 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roussopoulos:1986:PTD

[371] N. Roussopoulos and H. Kang. Principles and techniques in the design of ADMS. Computer, 19(12):19–25, December 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mark:1986:MM

[372] L. Mark and N. Roussopoulos. Metadata management. Computer, 19(12): 26–36, December 1986. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wiederhold:1986:VOD

[373] G. Wiederhold. Views, objects, and databases. Computer, 19(12):37–44, December 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Christodoulakis:1986:DPC

[374] S. Christodoulakis and C. Faloutsos. Design and performance considerations for an optical disk-based, multimedia object server. Computer, 19(12):45–56, December 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moldovan:1986:PMA

[375] D. I. Moldovan and J. A. B. Fortes. Partitioning and mapping algorithms into fixed-size systolic arrays. Computer, C-35(1):1–12, January 1986. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hwang:1987:CAA

[376] K. Hwang, J. Ghosh, and R. Chowkwanyun. Computer architectures for artificial intelligence processing. Computer, 20(1):19–27, January 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ungar:1987:WPS

[377] D. Ungar and D. Patterson. What price Smalltalk? Computer, 20(1):67– 74, January 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fahlman:1987:CAA

[378] S. E. Fahlman and Geoffrey E. Hinton. Connectionist architectures for artificial intelligence. Computer, 20(1):100– 109, January 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ketabchi:1987:MMC

[379] M. A. Ketabchi and V. Berzins. Modeling and managing CAD databases. Computer, 20(2):93–102, February 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berra:1987:CAS

[380] P. B. Berra, Soon M. Chung, and Nabil I. Hachem. Computer architecture for a surrogate file to a very large data/knowledge base. Computer, 20(3): 25–32, March 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pleszkun:1987:ALM

[381] A. R. Pleszkun and M. J. Thazhuthaveetil. The architecture of LISP machines. Computer, 20(3):35–44, March 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dongarra:1986:HDM

[382] J. Dongarra. How do the ‘minisupers’ stack up? Computer, 19 (3):93, 100, March 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://www.netlib.org/utk/people/
JackDongarra/PAPERS/How-Do-the-
Minisupers-Stack-Up.pdf.

Brooks:1987:NSB

[383] Frederick P. Brooks, Jr. No silver bullet: essence and accidents of software engineering. Computer, 20(4):10–19, April 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gevarter:1987:NEC

[384] W. B. Gevarter. The nature and evaluation of commercial expert system building tools. Computer, 20(5):24–41, May 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karp:1987:PP

[385] Alan H. Karp. Programming for parallelism. Computer, 20(5):43–57, May 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milutinovic:1987:ACS

[386] V. Milutinovic, D. Fura, W. Helbig, and J. Linn. Architecture/compiler synergism in GaAs computer systems. Computer, 20(5):72–93, May 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhuyan:1987:GEI

[387] L. N. Bhuyan. Guest Editor’s introduction: Interconnection networks for parallel and distributed processing. Computer, 20(6):9–13, June 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:1987:ASM

[388] Vijay P. Kumar and Sudhakar M. Reddy. Augmented shuffle-exchange multistage interconnection networks. Computer, 20(6):30–40, June 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Varma:1987:OCN

[389] A. Varma, A. A. Sawchuk, C. S. Raghavendra, and B. K. Jenkins. Optical crossbar networks. Computer, 20(6): 50–60, June 1987. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Fortes:1987:SAC

[390] J. A. B. Fortes and B. W. Wah. Systolic arrays — from concept to implementation. Computer, 20(7):12–17, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kung:1987:WAP

[391] S. Y. Kung, S. N. Lo, and J. N. Hwang.

Wavefront array processors — concept to implementation. Computer, 20(7): 18–34, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foulser:1987:SMA

[392] D. E. Foulser and R. Schreiber. The Saxpy Matrix-1: A general-purpose systolic computer. Computer, 20(7):35– 44, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drake:1987:SAS

[393] B. L. Drake, F. T. Luk, J. M. Speiser, and J. J. Symanski. SLAPP: A systolic linear algebra parallel processor. Computer, 20(7):45–50, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McCanny:1987:SSA

[394] J. V. McCanny and J. G. McWhirter.

Some systolic array developments in the United Kingdom. Computer, 20(7): 51–64, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abraham:1987:FTT

[395] J. A. Abraham, P. Banerjee, C. Chen, W. K. Fuchs, S. Y. Kuo, and A. L. N. Reddy. Fault tolerance techniques for systolic arrays. Computer, 20(7):65– 76, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Navarro:1987:PES

[396] J. J. Navarro, J. M. Llaberia, and M. Valero. Partitioning: An essential step in mapping algorithms into systolic array processors. Computer, 20(7): 77–90, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hein:1987:DGS

[397] C. E. Hein, R. M. Zieger, and J. A. Urbano. The design of a GaAs systolic array for an adaptive null steering beamforming controller. Computer, 20(7): 92–93, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nash:1987:SCS

[398] J. G. Nash, K. W. Przytula, and S. Hansen. The systolic/cellular system for signal processing. Computer, 20(7): 96–97, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:1987:ISA

[399] W. T. Lin, C. Y. Chin, and C. Y. Ho. Integrating systolic arrays into a supersystem. Computer, 20(7):100–101, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koren:1987:CID

[400] Israel Koren. The concept and implementation of data-driven processor arrays. Computer, 20(7):102–103, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhanu:1987:COR

[401] Bir Bhanu and Chih-Cheng Ho. CADbased 3D object representation for robot vision. Computer, 20(8):19–35, August 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roach:1987:CSB

[402] John W. Roach, Praveen K. Paripati, and Jeffrey S. Wright. A CAD system based on spherical dual representations. Computer, 20(8):37–44, August 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shneier:1987:PBV

[403] Michael O. Shneier, Ronald Lumia, and Martin Herman. Prediction-based vision for robot control. Computer, 20(8):46– 55, August 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horaud:1987:MSP

[404] Radu Horaud and Thomas Skordas.

Model-based strategy planning for recognizing partially occluded parts. Computer, 20(8):58–65, August 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gunnarsson:1987:CML

[405] Kristjan T. Gunnarsson and Friedrich B.

Prinz. CAD model-based localization of parts in manufacturing. Computer, 20 (8):66–74, August 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stevenson:1987:MPC

[406] Charles N. Stevenson. Model-based programming and control of robot manipulators. Computer, 20(8):76–84, August 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:FC

[407] Anonymous. [Front cover]. Computer, 20(9):c1–c2, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:TC

[408] Anonymous. Table of contents. Computer, 20(9):2–3, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Russo:1987:CSP

[409] R. Russo and M. Smith. Computer Society President’s message. Computer, 20 (9):6–7, September 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shriver:1987:ECM

[410] B. D. Shriver. Editor-in-Chief’s message. Computer, 20(9):10, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:FJC

[411] Anonymous. The 1987 Fall Joint Computer Conference Infomart — Dallas, Texas, October 25–29, 1987. Computer, 20(9):12–15, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conklin:1987:HIS

[412] Jeff Conklin. Hypertext: An introduction and survey. Computer, 20(9):17– 41, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boehm:1987:ISP

[413] Boehm. Improving software productivity. Computer, 20(9):43–57, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gimarc:1987:SRP

[414] C. E. Gimarc and V. M. Milutinovic. A survey of RISC processors and computers of the mid-1980s. Computer, 20(9): 59–69, September 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flynn:1987:NCM

[415] M. J. Flynn, C. L. Mitchell, and J. M.

Mulder. And now a case for more complex instruction sets. Computer, 20(9): 71–83, September 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoare:1987:OSF

[416] C. A. R. Hoare. An overview of some formal methods for program design. Computer, 20(9):85–91, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:S

[417] Anonymous. Standards. Computer, 20 (9):94–97, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:U

[418] Anonymous. Update. Computer, 20(9): 98–99, September 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:CSN

[419] Anonymous. Computer Society news. Computer, 20(9):99–102, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:NPR

[420] Anonymous. New product reviews. Computer, 20(9):103–106, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:NP

[421] Anonymous. New products. Computer, 20(9):107–113, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:IA

[422] Anonymous. IC announcements. Computer, 20(9):114, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:MA

[423] Anonymous. Microsystem announcements. Computer, 20(9):115, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:IIC

[424] Anonymous. IEEE International Conference on Computer-Aided Design. Computer, 20(9):116–117, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:Ca

[425] Anonymous. Conferences. Computer, 20 (9):118–119, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:Cb

[426] Anonymous. Calendar. Computer, 20 (9):119–120, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:CP

[427] Anonymous. Call for papers. Computer, 20(9):121–122, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:CO

[428] Anonymous. Career opportunities.

Computer, 20(9):122–123, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:BR

[429] Anonymous. Book reviews. Computer, 20(9):124–126, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:NL

[430] Anonymous. New literature. Computer, 20(9):127, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:CSM

[431] Anonymous. The Computer Society: a member society of the Institute of Electrical and Electronics Engineers, Inc. Computer, 20(9):129, September 1987.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stone:1987:PQL

[432] Harold S. Stone. Parallel querying of large databases: A case study. Computer, 20(10):11–21, October 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berra:1987:OTD

[433] P. B. Berra and N. B. Troullinos. Optical techniques and data/knowledge base machines. Computer, 20(10):59– 70, October 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Henderson:1987:IDP

[434] P. B. Henderson and D. Notkin. Integrated design and programming environments. Computer, 20(11):12–16, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dart:1987:SDE

[435] S. A. Dart, R. J. Ellison, P. H. Feiler, and N. A. Habermann. Software development environments. Computer, 20 (11):18–28, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reps:1987:LPP

[436] T. Reps and T. Teitelbaum. Language processing in program editors. Computer, 20(11):29–40, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goguen:1987:FPE

[437] Joseph Goguen and Mark Moriconi.

Formalization in programming environments. Computer, 20(11):55–64, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carle:1987:PES

[438] Alan Carle, Keith D. Cooper, Robert T.

Hood, Ken Kennedy, Linda Torczon, and Scott K. Warren. A practical environment for scientific programming. Computer, 20(11):75–89, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Denning:1988:CDP

[439] P. J. Denning, D. E. Comer, D. Gries, M. C. Mulder, A. Tucker, A. J. Turner, and P. R. Young. Computing as a discipline: Preliminary report of the ACM Task Force on the Core of Computer Science. Computer, 20(1):41, February 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also final report [506].

Chen:1988:MRD

[440] P. Chen and M. A. Harrison. Multiple representation document development. Computer, 21(1):15–31, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1988:SDE

[441] J. Johnson and R. J. Beach. Styles in document editing systems. Computer, 21(1):32–47, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schlichter:1988:FAP

[442] J. H. Schlichter and L. J. Miller. FolioPub: A publication management system. Computer, 21(1):61–69, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marchionini:1988:FFV

[443] G. Marchionini and B. Shneiderman. Finding facts vs. browsing knowledge in hypertext systems. Computer, 21(1):70– 80, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yankelovich:1988:ICC

[444] Nicole Yankelovich, Bernard Hann, Norman Meyrowitz, and S. M. Drucker. Intermedia: The concept and the construction of a seamless information environment. Computer, 21(1):81–97, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dubois:1988:SCE

[445] Michel Dubois, Christoph Scheurich, and Fayé A. Briggs. Synchronization, coherence, and event ordering in multiprocessors. Computer, 21(2):9–21, February 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ousterhout:1988:SNO

[446] John K. Ousterhout, A. R. Cherenson, Fred Douglis, Michael N. Nelson, and Brent B. Welch. The Sprite network operating system. Computer, 21(2):23– 36, February 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bernstein:1988:SAF

[447] P. A. Bernstein. Sequoia: A faulttolerant tightly coupled multiprocessor for transaction processing. Computer, 21 (2):37–45, February 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lang:1988:OPA

[448] Gordon R. Lang, Moez Dharssi, Fred M.

Longstaff, Philip S. Longstaff, Peter A. S. Metford, and Malcolm T. Rimmer. An optimum parallel architecture for high-speed real-time digital signal processing. Computer, 21(2):47– 57, February 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moffett:1988:SAC

[449] Jonathan D. Moffett and Morris S. Sloman. The source of authority for commercial access control. Computer, 21 (2):59–69, February 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Widrow:1988:NNA

[450] B. Widrow and R. Winter. Neural nets for adaptive filtering and adaptive pattern recognition. Computer, 21(3):25– 39, March 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Graf:1988:VIN

[451] Hans P. Graf, Lawrence D. Jackel, and Wayne E. Hubbard. VLSI implementation of a neural network model. Computer, 21(3):41–49, March 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hutchinson:1988:CMU

[452] James Hutchinson, Christof Koch, Jin Luo, and Carver Mead. Computing motion using analog and binary resistive networks. Computer, 21(3):52– 63, March 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fukushima:1988:NNF

[453] Kunihiko Fukushima. Neural network for visual pattern recognition. Computer, 21(3):65–75, March 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carpenter:1988:AAP

[454] Gail A. Carpenter and Stephen Grossberg. The ART of adaptive pattern recognition by a self-organizing neural network. Computer, 21(3):77–88, March 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feldmann:1988:CSN

[455] Jerome A. Feldmann, Mark A. Fanty, and Nigel H. Goddard. Computing with structured neural networks. Computer, 21(3):91–103, March 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Butler:1988:MLG

[456] J. T. Butler. Multiple-valued logic: Guest Editor’s introduction and bibliography. Computer, 21(4):13–??, April 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1988:MLA

[457] K. C. Smith. Multiple-valued logic: A tutorial and appreciation. Computer, 21(4):17–27, April 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Etiemble:1988:CBM

[458] D. Etiemble and M. Israel. Comparison of binary and multivalued ICs according to VLSI criteria. Computer, 21(4): 28–42, April 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kameyama:1988:MCM

[459] M. Kameyama, S. Kawahito, and T. Higuchi. A multiplier chip with multiple-valued bidirectional currentmode logic circuits. Computer, 21(4): 43–57, April 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Butler:1988:MCC

[460] J. T. Butler and H. G. Kerkhoff.

Multiple-valued CCD circuits. Computer, 21(4):58–70, April 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sasao:1988:MLO

[461] T. Sasao. Multiple-valued logic and optimization of programmable logic arrays. Computer, 21(4):71–82, April 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nichols:1988:MMS

[462] Kathleen M. Nichols and John T. Edmark. Modeling multicomputer systems with PARET. Computer, 21(5): 39–48, May 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bershad:1988:RCF

[463] Brian N. Bershad and Henry M. Levy. A remote computation facility for a heterogeneous environment. Computer, 21(5): 50–60, May 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boehm:1988:SMS

[464] Barry W. Boehm. A spiral model of software development and enhancement. Computer, 21(5):61–72, May 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gait:1988:OFC

[465] Jason Gait. The optical file cabinet: A random-access file system for writeonce optical disks. Computer, 21(6): 11–22, June 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mills:1988:SRV

[466] H. D. Mills. Stepwise refinement and verification in box-structured systems. Computer, 21(6):23–37, June 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rayfield:1988:SAS

[467] J. T. Rayfield and H. F. Silverman. System and application software for the Armstrong multiprocessor. Computer, 21(6):38–52, June 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Potter:1988:TSH

[468] Walter D. Potter and Robert P. Trueblood. Traditional, semantic, and hypersemantic approaches to data modeling.

Computer, 21(6):53–63, June 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:1988:PVF

[469] Bowen Liu and Nelson Strother. Programming in VS Fortran on the IBM 3090 for maximum vector performance. Computer, 21(6):65–77, June 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Camurati:1988:FVH

[470] P. Camurati and P. Prinetto. Formal verification of hardware correctness: Introduction and survey of current research. Computer, 21(7):8–20, July 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hooton:1988:EPD

[471] A. R. Hooton, U. Agueero, and S. Dasgupta. An exercise in plausibility-driven design. Computer, 21(7):21–33, July 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Steenkiste:1988:LRP

[472] P. Steenkiste and J. Hennessy. Lisp on a reduced-introduction-set processor: Characterization and optimization. Computer, 21(7):34–46, July 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athas:1988:MMP

[473] William C. Athas and Charles L. Seitz.

Multicomputers: Message-passing concurrent computers. Computer, 21(8):9– 24, August 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tucker:1988:AAC

[474] Lewis W. Tucker and George G. Robertson. Architecture and applications of the Connection Machine. Computer, 21(8): 26–38, August 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rich:1988:APM

[475] Charles Rich and Richard C. Waters. Automatic programming: Myths and prospects. Computer, 21(8):40– 51, August 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gries:1988:TS

[476] David Gries and Dorothy Marsh. The 1986–1987 Taulbee Survey. Computer, 21(8):53–63, August 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Diede:1988:TGS

[477] Tom Diede, Carl F. Hagenmaier, Glen S. Miranker, Jonathan J. Rubinstein, and William S. Worley, Jr. The Titan Graphics Supercomputer architecture. Computer, 21(9):13–28, 30, September 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolcott:1988:HCS

[478] Peter Wolcott and Seymour E. Goodman. High-speed computers of the Soviet Union. Computer, 21(9):32–41, September 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leung:1988:FCE

[479] K. S. Leung and W. Lam. Fuzzy concepts in expert systems. Computer, 21 (9):43–56, September 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1988:NCS

[480] W. Myers and M. Potes. National Conference on Strategic Management of R&D. Computer, 21(9):58–67, September 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stankovic:1988:MAR

[481] J. A. Stankovic. Misconceptions about real-time computing — A serious problem for next-generation systems. Computer, 21(10):10–19, October 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Comment 1 by schlenk, Mon Dec 12 13:13:04 1988 An interesting outlook on real-time systems that are integrated into a programming environment is given. The author shows why ADA tasks are not sufficient (a high priority task has to wait for a lower priority task) and suggests an integrated approach: Time management on the language level, schedulability check and support for distributed programs.

Doran:1988:AMA

[482] R. W. Doran. Amdahl multiple-domain architecture. Computer, 21(10):20– 28, October 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koren:1988:DVA

[483] Israel Koren, Bilha Mendelson, Irit Peled, and Gabriel M. Silberman. A data-driven VLSI array for arbitrary algorithms. Computer, 21(10):30–43, October 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1988:HDS

[484] Ted G. Lewis and Curtis R. Cook. Hashing for dynamic and static internal tables. Computer, 21(10):45–56, October 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). The authors survey the classical hashing function approach to information retrieval and show how general hashing techniques exchange speed for memory. It is a tutorial paper that covers, among other topics, dynamic and static hash tables, perfect hashing, and minimal perfect hashing.

Lie:1988:AML

[485] Y.-C. D. Lie. Appendum to ‘Multiplevalued logic: a tutorial and appreciation’ (April 17–27, 1988). Computer, 21(11): 9, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rich:1988:PAR

[486] C. Rich and R. C. Waters. The Programmer’s Apprentice: a research overview. Computer, 21(11):10–25, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stenstrom:1988:RCS

[487] P. Stenström. Reducing contention in shared-memory multiprocessors. Computer, 21(11):26–37, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Britcher:1988:UII

[488] R. N. Britcher. Using inspections to investigate program correctness. Computer, 21(11):38–44, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Skillicorn:1988:TCA

[489] David B. Skillicorn. A taxonomy for computer architectures. Computer, 21 (11):46–57, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Webster:1988:MDI

[490] D. E. Webster. Mapping the design information representation terrain. Computer, 21(12):8–23, December 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hill:1988:CDC

[491] M. D. Hill. A case for direct-mapped caches. Computer, 21(12):25–40, December 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaffner:1988:SSD

[492] Stuart C. Schaffner and Martha Borkan.

Segue: Support for distributed graphical interfaces. Computer, 21(12):42– 55, December 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Detlefs:1988:ISR

[493] David L. Detlefs, Maurice P. Herlihy, and Jeannette M. Wing. Inheritance of synchronization and recovery properties in Avalon/C++. Computer, 21(12):57– 69, December 1988. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1988:SS

[494] Anonymous. Software for supercomputers. Computer, 21(12):70–74, December 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rau:1989:CDS

[495] B. Ramakrishna Rau, D. W. L. Yen, Wei Yen, and R. A. Towle. The Cydra 5 departmental supercomputer: design philosophies, decisions, and tradeoffs. Computer, 22(1):12–35, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1989:EDC

[496] T. Jones. Engineering design of the Convex C2. Computer, 22(1):36–44, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

August:1989:CXB

[497] M. C. August, G. M. Brost, C. C. Hsiung, and A. J. Schiffleger. Cray X-MP: The birth of a supercomputer. Computer, 22(1):45–52, January 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cragon:1989:TAS

[498] H. G. Cragon and W. J. Watson. The TI Advanced Scientific Computer. Computer, 22(1):55–64, January 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maytal:1989:DCG

[499] B. Maytal, S. Iacobovici, D. B. Alpert, D. Biran, J. Levy, and S. Y. Tov. Design considerations for a general-purpose microprocessor. Computer, 22(1):66– 76, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1989:PA

[500] R. B. Lee. Precision architecture. Computer, 22(1):78–91, January 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1989:SPS

[501] W. Myers. Software pivotal to strategic defense. Computer, 22(1):92–97, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Linton:1989:CUI

[502] Mark A. Linton, John M. Vissides, and Paul R. Calder. Composing user interfaces with InterViews. Computer, 22 (2):8–22, February 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). From Computing Reviews: “InterViews objects are combined through a model similar to the boxes and glues in Knuth’s TEX, which allows interface objects to be combined with little concern for undesired interactions.”.

Bhuyan:1989:PMI

[503] L. N. Bhuyan, Qing Yang, and D. P.

Agrawal. Performance of multiprocessor interconnection networks. Computer, 22 (2):25–37, February 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Richardson:1989:COA

[504] Stephen Richardson and Mahadevan Ganapathi. Code optimization across procedures. Computer, 22(2):42–50, February 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ganapathi:1989:IAC

[505] M. Ganapathi and G. O. Mendal. Issues in Ada compiler technology. Computer, 22(2):52–60, February 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Denning:1989:CD

[506] P. J. Denning, D. E. Comer, D. Gries, M. C. Mulder, A. Tucker, A. J. Turner, and P. R. Young. Computing as a discipline. Computer, 22(2):63–70, February 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also preliminary report [439].

Gries:1989:SRD

[507] D. Gries, T. Walker, and P. R. Young. 1988 Snowbird report: a discipline matures (computer science education). Computer, 22(2):72–75, February 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilfong:1989:RA

[508] G. T. Wilfong. Robotics and automation. Computer, 22(3):6–8, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sharir:1989:AMP

[509] M. Sharir. Algorithmic motion planning in robotics. Computer, 22(3):9– 19, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lozano-Perez:1989:TPP

[510] T. Lozano-Perez, J. L. Jones, E. Mazer, and P. A. O’Donnell. Task-level planning of pick-and-place robot motions. Computer, 22(3):21–29, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoffmann:1989:PAR

[511] C. M. Hoffmann. The problems of accuracy and robustness in geometric computation. Computer, 22(3):31–39, 41, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cox:1989:EHR

[512] I. J. Cox and N. H. Gehani. Exception handling in robotics. Computer, 22(3): 43–49, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Allen:1989:CRR

[513] P. Allen, P. Michelman, and K. Roberts.

Current research in robotics and automation — an intelligent grasping system. Computer, 22(3):50–52, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eicker:1989:CRR

[514] P. J. Eicker and D. R. Strip. Current research in robotics and automation — automated planning and programming for robotic batch mechanical assembly. Computer, 22(3):53–54, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hopcroft:1989:CRR

[515] J. E. Hopcroft. Current research in robotics and automation — electronic prototyping. Computer, 22(3): 55–57, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bekey:1989:CRR

[516] G. A. Bekey. Current research in robotics and automation-robotics research at USC. Computer, 22(3):58– 60, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keefe:1989:SQS

[517] T. F. Keefe, M. B. Thuraisingham, and W. T. Tsai. Secure queryprocessing strategies. Computer, 22(3): 63–70, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhattacharya:1989:HTG

[518] Debashis Bhattacharya, Brian T. Murray, and John P. Hayes. High-level test generation for VLSI. Computer, 22(4): 16–24, April 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Renous:1989:WWT

[519] R. Renous, G. M. Silberman, and I. Spillinger. Whistle: A workbench for test development of library-based designs. Computer, 22(4):27–41, April 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheng:1989:GAT

[520] W.-T. Cheng and T. J. Chakraborty.

Gentest: An automatic test-generation system for sequential circuits. Computer, 22(4):43–49, April 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Heap:1989:GSC

[521] M. A. Heap and W. A. Rogers. Generating single-stuck-fault coverage from a collapsed-fault set. Computer, 22(4): 51–57, April 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hallenbeck:1989:TEA

[522] J. J. Hallenbeck, J. R. Cybrynski, N. Kanopoulos, T. Markas, and N. Vasanthavada. The Test Engineer’s Assistant: A support environment for hardware design for testability. Computer, 22(4):59–68, April 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tanik:1989:RPS

[523] M. M. Tanik and R. T. Yeh. Rapid prototyping in software development. Computer, 22(5):9–12, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Luqi:1989:SET

[524] Luqi. Software evolution through rapid prototyping. Computer, 22(5): 13–25, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:1989:OVC

[525] R. Gupta, W. H. Cheng, R. Gupta, I. Hardonag, and M. A. Breuer. An object-oriented VLSI CAD framework. Computer, 22(5):28–37, May 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jordan:1989:SSC

[526] P. W. Jordan, K. S. Keller, R. W. Tucker, and D. Vogel. Software storming: Combining rapid prototyping and knowledge engineering. Computer, 22 (5):39–48, May 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1989:PSU

[527] T. G. Lewis, F. Handloser, III, S. Bose, and S. Yang. Prototypes from Standard User Interface Management Systems. Computer, 22(5):51–60, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Duke:1989:RPF

[528] E. L. Duke, R. W. Brumbaugh, and J. D. Disbrow. A rapid prototyping facility for flight research in advanced systems concepts. Computer, 22(5):61–66, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cieslak:1989:PNP

[529] R. Cieslak, A. Fawaz, S. Sachs, P. Varaiya, J. Walrand, and A. Li. The programmable network prototyping system. Computer, 22(5):67–76, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McKeeman:1989:GTA

[530] W. M. McKeeman. Graduation talk at Wang Institute. Computer, 22(5): 78–80, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Iyengar:1989:AIM

[531] S. S. Iyengar and R. L. Kashyap. Autonomous intelligent machines. Computer, 22(6):14–17, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bares:1989:AAR

[532] John Bares, Martial Hebert, Takeo Kanade, Eric Krotkov, Tom Mitchell, Reid Simmons, and William Whittaker. Ambler — an autonomous rover for planetary exploration. Computer, 22(6): 18–26, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weisbin:1989:AMR

[533] C. R. Weisbin, G. de Saussure, J. R.

Einstein, F. G. Pin, and E. Heer. Autonomous mobile robot navigation and learning. Computer, 22(6):29–35, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rao:1989:AFL

[534] N. S. V. Rao. Algorithmic framework for learned robot navigation in unknown terrains. Computer, 22(6):37–43, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Elfes:1989:UOG

[535] A. Elfes. Using occupancy grids for mobile robot perception and navigation.

Computer, 22(6):46–57, June 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bihari:1989:CAS

[536] T. E. Bihari, T. M. Walliser, and M. R. Patterson. Controlling the adaptive suspension vehicle. Computer, 22(6): 59–65, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barhen:1989:NLC

[537] Jacob Barhen, Sandeep Gulati, and Michail Zak. Neural learning of constrained nonlinear transformations. Computer, 22(6):67–76, June 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1989:PRK

[538] I. Lee, R. B. King, and R. P. Paul. A predictable real-time kernel for distributed multisensor systems. Computer, 22(6): 78–83, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roth-Tabak:1989:BEM

[539] Y. Roth-Tabak and R. Jain. Building an environment model using depth information. Computer, 22(6):85–90, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trivedi:1989:VSR

[540] M. M. Trivedi, C. X. Chen, and S. B. Marapane. A vision system for robotic inspection and manipulation. Computer, 22(6):91–97, June 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Preston:1989:ACB

[541] K. Preston, Jr. The Abingdon Cross benchmark survey. Computer, 22(7): 9–18, July 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1989:DIS

[542] J. E. Smith. Dynamic instruction scheduling and the Astronautics ZS-1. Computer, 22(7):21–35, July 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Biggerstaff:1989:DRM

[543] T. J. Biggerstaff. Design recovery for maintenance and reuse. Computer, 22 (7):36–49, July 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chisvin:1989:CAM

[544] L. Chisvin and R. J. Duckworth.

Content-addressable and associative memory: Alternatives to the ubiquitous RAM. Computer, 22(7):51–64, July 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dinning:1989:SSM

[545] A. Dinning. A survey of synchronization methods for parallel computers. Computer, 22(7):66–77, July 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFanti:1989:VES

[546] T. A. DeFanti, M. D. Brown, and B. H.

McCormick. Visualization: Expanding scientific and engineering research opportunities. Computer, 22(8):12–16, 22– 25, August 1989. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Helman:1989:RDV

[547] J. Helman and L. Hesselink. Representation and display of vector field topology in fluid flow data sets. Computer, 22(8): 27–36, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Long:1989:ARD

[548] M. B. Long, K. Lyons, and J. K. Lam. Acquisition and representation of 2D and 3D data from turbulent flows and flames. Computer, 22(8):39–45, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuchs:1989:IVM

[549] Henry Fuchs, Marc Levoy, and Stephen M. Pizer. Interactive visualization of 3-D medical data. Computer, 22(8):46–51, August 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hibbard:1989:VLD

[550] W. Hibbard and D. Santek. Visualizing large data sets in the Earth sciences. Computer, 22(8):53–59, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kluksdahl:1989:RVS

[551] N. C. Kluksdahl, A. M. Kriman, and D. K. Ferry. The role of visualization in the simulation of quantum electronic transport in semiconductors. Computer, 22(8):60–66, August 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Phillips:1989:DVA

[552] R. L. Phillips. Distributed visualization at Los Alamos National Laboratory. Computer, 22(8):70–77, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cabral:1989:VTA

[553] B. Cabral and C. L. Hunter. Visualization tools at Lawrence Livermore National Laboratory. Computer, 22(8):77– 84, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haber:1989:SVR

[554] R. B. Haber. Scientific visualization and the Rivers Project at the National Center for Supercomputing Applications. Computer, 22(8):84–89, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bancroft:1989:SVC

[555] G. V. Bancroft, T. Plessel, F. Merritt, P. P. Walataka, and V. Watson. Scientific visualization in computational aerodynamics at NASA Ames Research Center. Computer, 22(8):89–95, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rosenblum:1989:VED

[556] L. J. Rosenblum. Visualization of experimental data at the Naval Research Laboratory. Computer, 22(8):95– 101, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pappas:1989:TIP

[557] T. L. Pappas. TEXnology on the IBM PC. Computer, 22(8):111–120, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1989:XSR

[558] J. Johnson, T. L. Roberts, W. Verplank, D. C. Smith, C. H. Irby, M. Beard, and K. Mackey. The Xerox Star: a retrospective. Computer, 22(9):11–26, 28– 29, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheng:1989:VPC

[559] H. Cheng. Vector pipelining, chaining, and speed on the IBM 3090 and Cray X-MP. Computer, 22(9):31–42, 44, 46, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lieberherr:1989:AGS

[560] Karl J. Lieberherr and Ian M. Holland. Assuring good style for objectoriented programs. Computer, 22(9):38– 44, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Artsy:1989:DPM

[561] Yeshayahu Artsy and Raphael Finkel. Designing a process migration facility: The Charlotte experience. Computer, 22(9):47–56, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Review: Computing Reviews, Vol. 31, No. 8, August 1990.

Ford:1989:MSE

[562] G. A. Ford and N. E. Gibbs. A Master of Software Engineering Curriculum: Recommendations from the Software Engineering Institute. Computer, 22(9):59– 71, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ambler:1989:IVT

[563] A. L. Ambler and M. M. Burnett. Influence of visual technology on the evolution of language environments. Computer, 22(10):9–22, October 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roman:1989:DAV

[564] G.-C. Roman and K. C. Cox. A declarative approach to visualizing concurrent computations. Computer, 22(10):25– 36, October 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lehr:1989:VPD

[565] Ted Lehr, Zary Segall, Dalibor F. Vrsalovic, Eddie Caplan, Alan L. Chung, and Charles E. Fineman. Visualizing performance debugging. Computer, 22(10):38– 51, October 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kramer:1989:GCP

[566] Jeffrey Kramer, J. Magee, and K. Ng.

Graphical configuration programming. Computer, 22(10):53–58, 62–65, October 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gu:1989:SAV

[567] J. Gu and K. F. Smith. A structured approach for VLSI circuit design. Computer, 22(11):9–22, November 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hac:1989:SON

[568] A. Hac and H. B. Mutlu. Synchronous optical network and broadband ISDN protocols. Computer, 22(11):26–34, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singhal:1989:DDD

[569] M. Singhal. Deadlock detection in distributed systems. Computer, 22(11):37– 48, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gries:1989:TSR

[570] D. Gries and D. Marsh. The 1987–88 Taulbee Survey Report. Computer, 22 (11):49–56, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1989:CSP

[571] Anonymous. The computer spectrum: a perspective on the evolution of computing. Computer, 22(11):57–63, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1989:SHU

[572] Anonymous. Supercomputer hardware: an update of the 1983 report’s summary and tables. Computer, 22(11):63– 68, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grosky:1989:GEI

[573] W. I. Grosky and R. Mehrotra. Guest Editors’ introduction: Image database management. Computer, 22(12):7–8, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kasturi:1989:MDP

[574] R. Kasturi, R. Fernandez, M. L. Amlani, and W.-C. Feng. Map data processing in Geographic Information Systems. Computer, 22(12):10–21, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brolio:1989:IDS

[575] J. Brolio, B. A. Draper, J. R. Beveridge, and A. R. Hanson. ISR: A database for symbolic processing in computer vision. Computer, 22(12):22–30, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jagadish:1989:OMI

[576] H. V. Jagadish and L. O’Gorman.

An object model for image recognition. Computer, 22(12):33–41, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goodman:1989:KBC

[577] Andrew M. Goodman, Robert M. Haralick, and Linda G. Shapiro. Knowledgebased computer vision — integrated programming language and data management system design. Computer, 22(12): 43–54, December 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pizano:1989:SSI

[578] A. Pizano, A. Klinger, and A. Cardenas. Specification of spatial integrity constraints in pictorial databases. Computer, 22(12):59–71, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ziv:1977:UAS

[579] Jacob Ziv and Abraham Lempel. A universal algorithm for sequential data compression. IEEE Transactions on Information Theory, IT-23(3):337–343, May 1977. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Ziv:1978:CIS

[580] J. Ziv and A. Lempel. Compression of individual sequences via variable rate coding. IEEE Transactions on Information Theory, IT-24(5):530–536, September 1978. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Crocker:1995:PPN

[581] Lee Daniel Crocker. PNG: The Portable Network Graphic format. Dr. Dobb’s Journal of Software Tools, 20(7):36, 38, 40, 42, 44, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Roelofs:1999:PDG

[582] Greg Roelofs. PNG: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-542-4. xix + 327 + 4 pp. LCCN T385 .R58 1999. US$32.95. URL http://www.

libpng.org/pub/png/pngbook.html;
 http://www.oreilly.com/catalog/
pngdefg.

Salomon:2000:DCC

[583] David Salomon. Data Compression:

The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-95045-1. xvi + 823 pp. LCCN QA76.9.D33 S25 2000. US$39.95. URL http://

www.ecs.csun.edu/~dxs/DCadvertis/
Dcomp2Ad.html.